### v1

# JUREX-4E: Juridical Expert-Annotated Four-Element Knowledge Base for Legal Reasoning

Abstract
The Four-Element Theory is a fundamental framework in criminal law, defining the constitution of crime through four dimensions: Subject, Object, Subjective aspect, and Objective aspect. This theory is widely referenced in legal reasoning, and many Large Language Models (LLMs) attempt to incorporate it when handling legal tasks. However, current approaches rely on LLMs’ internal knowledge to incorporate this theory, often lacking completeness and representativeness. To address this limitation, we introduce JUREX-4E, an expert-annotated knowledge base covering 155 criminal charges. It is structured through a progressive hierarchical annotation framework that prioritizes legal source validity and employs diverse legal interpretation methods to ensure comprehensiveness and authority. We evaluate JUREX-4E on the Similar Charge Distinction task and apply it to Legal Case Retrieval, demonstrating its effectiveness in improving LLM performance. Experimental results validate the high quality of JUREX-4E and its substantial impact on downstream legal tasks, underscoring its potential for advancing legal AI applications. Code: https://github.com/THUlawtech/JUREX
JUREX-4E: Juridical Expert-Annotated Four-Element Knowledge Base for Legal Reasoning
[…]
1 Introduction
In legal AI tasks, enhancing the accuracy and interpretability of Large Language Models (LLMs) in the legal domain often requires the incorporation of legal theories as a support Jiang and Yang (2023); Servantez et al. (2024); Yuan et al. (2024); Deng et al. (2023). One important theory is the Four-Element Theory of Crime Constitution in Chinese criminal law Liang (2017). This theory deconstructs criminal conduct into four elements: Subject, Object, Subjective aspect, and Objective aspect, providing clear standards for judicial authorities to determine criminal behavior and helping to prevent the abuse of penal power.
However, most current approaches do not provide additional knowledge but rather rely on the LLM’s internal knowledge to incorporate the Four-Element Theory. A common method is to guide LLMs in mimicking expert reasoning processes. For example, designing four separate prompts to guide the LLM outputs in the form of four elementsDeng et al. (2023).
These methods assume that the model has a solid grasp of the Four-Element Theory, which has not yet been verified. We had LLMs generate the four elements of several complicated crimes in Chinese judicial practiceOuyang et al. (1999), and then asked legal experts to score them. We found that, although LLMs can generate formally standardized and relatively accurate legal descriptions when provided with legal theoretical frameworks and references, the model still underperformed in terms of completeness and representativeness. This shortcoming could affect the accuracy and soundness of subsequent reasoning.
To help LLMs better utilize the Four-Element Theory in legal tasks, we propose JUREX-4E: JURidical EXpert-annotated 4-Element knowledge base for legal reasoning. This knowledge base is annotated using a progressive hierarchy: Article → Judicial Interpretations → Guiding Cases → Academic Discourses, which is built upon the pyramid structure of legal source validity. It incorporates various legal interpretation methods, including textual, systematic, sociological, and purposive interpretations. The knowledge base covers the four elements of 155 high frequency charges, annotated by legal experts over a period of seven months. Each crime’s four elements are described in an average of 472.5 words.
To assess the quality of the annotations, we sampled several crimes for human evaluation. The expert annotations achieved an average score of 4.60 on a 5-point scale, while the LLM-generated four elements scored only 3.96, indicating that the expert annotations were of higher quality. To further evaluate the annotations objectively and comprehensively, a direct way is to judge whether different charges can be distinguished according to the four-element definition of crime constitution. Therefore, we introduced the Similar Charge Distinction task Liu et al. (2021). For each case, we provided the four elements of the candidate confused charges and combined them with the case facts as model input. The experimental results showed that injecting expert annotations helped the model better differentiate between similar charges, improving performance with a 0.65 increase in average accuracy and a 0.70 increase in average F1-score, underscoring the superior quality and reliability of expert annotations compared to those generated by the LLM.
We also applied the expert annotations in a specific legal task: Legal Case Retrieval. It is an important step in the practice of analyzing cases and making judgments, requiring the precise application of the four-element theory to compare the criminal composition of cases. We designed a simple retrieval framework guided by expert knowledge, in which the charge’s four elements was used to generate four-element descriptions for both the query and candidate cases, and then match similar cases based on their vector similarity. Experiments demonstrated that incorporating expert-annotated four elements improved retrieval performance, as the model became better at focusing on the legal features and key details.
Our contributions are as follows:
  1. (1)
    We verify that LLMs have gaps in understanding the legal theory, highlighting the inadequacy of relying solely on LLM-driven reasoning for legal AI tasks.
  2. (2)
    We built the JUREX-4E knowledge base, which is the first to incorporate the pyramid structure of legal source validity and covers the four elements of 155 criminal charges under Chinese Criminal Law.
  3. (3)
    We demonstrated the significance of incorporating criminal composition elements in the Similar Charge Distinction task and proved the superior quality of the expert-annotated four-element knowledge base.
  4. (4)
    We applied JUREX-4E to the Legal Case Retrieval task, found that they do indeed contribute to downstream tasks.
2 Related Work
In legal AI, much work has introduced legal theories to enhance reasoning and improve model accuracy and interpretability. For example, legal syllogism prompting (LoT)Jiang and Yang (2023) teaches LLMs for legal judgment prediction by instructing legal syllogism, Chain of LogicServantez et al. (2024) guides models in reasoning about compositional rules by decomposing logical statements based on the IRAC (Issue, Rule, Application, Conclusion) paradigm. Among these, the Four-Elements Theory (FET) of Crime Constitution is a widely adopted frameworkYuan et al. (2024); Deng et al. (2023).
The Four-Element Theory is one of the most widely recognized criminal theories in Chinese judicial practice Liang (2017). It specifies four essential elements that must be satisfied to establish criminal liability: Subject, Object, Subjective aspect, and Objective aspect. For example, the four elements of the Crime of Affray can be briefly summarized as follows:
(1) Subject: Principal organizers and other active participants who have reached the age of criminal responsibility. (2) Object: Public order. (3) Objective Aspect: The act of assembling brawl, engaging in a brawl, resulting in the following consequences of serious injury. (4) Subjective Aspect: Direct intent, where the person knowingly and willfully engages in organizing or participating in the act of assembling brawl.
Before discussing the Four-Element Theory (FET), it is necessary to briefly compare it with another key theory in Chinese criminal law, the Hierarchical Theory of Crime ConstitutionZhou (2017b); Zhang (2010), and the main distinction between these theories lies in whether a hierarchical structure is considered, with ongoing debates in practiceGao (2009); Chen (2010, 2017); Zhou (2017a). We chose FET as our foundational template for following reasons: 1) its dominance in Chinese judicial practice aligns with real-world criminal judgments; (2) its clear distinction between objective aspects and subjective intent offers direct reasoning checkpoints compared to the Three-Tier Theory; (3) its four-element annotation is flexible and can be adapted to the Three-Tier Theory by prioritizing objective analysis before subjective evaluationLi (2006); Zhang (2017).
Recent approaches have leveraged the FET framework to model expert reasoning. For example, breaking down legal rules into FET-aligned components using automated planning techniques Yuan et al. (2024). Employing model-generated four-element structures as minor premises in legal judgment analysis Deng et al. (2023). While these methods have demonstrated improved performance on downstream tasks, they generally assume that the LLMs inherently understand the FET, without systematically validating this assumption. Notably, prior research on criminal charge prediction An et al. (2022) suggests that the models may misinterpret key legal concepts and may not be sensitive enough to the subtle differences in fact descriptions of confusing charges, highlighting the need to incorporate expert annotations to support LLM reasoning.
3 Dataset Construction
3.1 Hierarchical Legal Interpretation System
Figure 1: Hierarchical Legal Interpretation System base on legal source validity. The legal sources follow a hierarchical order of validity. Thick arrows indicate the primary level where a particular interpretive method is applied, while dashed arrows represent its supplementary use at that level.
Annotating the four elements of crime constitution is essentially a process of legal interpretation, which can be analyzed in two aspects:
(1) What law is being interpreted. This involves identifying the sources of law, including statutory provisions corresponding to a specific charge, their associated judicial interpretations, case precedents, and academic discourses. In legal studies, these sources are categorized based on their legal validity into formal sources (which carry legal forces in judgments) and informal sources (which serve as references without legal forces)Pound (1925); Watson (1982); Pound (1932). Articles and judicial interpretations are considered formal sources, whereas case precedents and academic discourses are regarded as informal sources under the Chinese legal systemZhang and Zhou (2007).
(2) How the law is interpreted. This pertains to legal interpretation methods, including literal interpretation, systematic interpretation, purposive interpretation, etc. These methods follow a hierarchical order in legal reasoningSutherland (1891); Kim and Division (2008); Eig Larry (2014). Legal interpretation should begin with literal interpretation (textual analysis). If the intended meaning cannot be clearly derived from the article alone, systematic interpretation and purpose interpretation should be applied first. If ambiguity remains, historical interpretation and comparative law interpretation may be used to further clarify the legal meaning. The specific definition is in AppendixB.
Based on these principles, our annotation follows a pyramid structure of Hierarchical Legal Interpretation System base on legal source validity. As shown in Figure 1, the system is divided into two parts: Legal Source and Legal Interpretation Methods. The main structure of legal source follows a hierarchical order of validity: Article → Judicial Interpretations → Guiding Cases → Academic Discourses, where various legal interpretation methods are applied across different levels. Thick arrows indicate the primary level at which a particular method is used, while thin arrows denote the cross applications.
3.2 Hierarchical Annotation Path of Legal Sources
Our Annotators are experts who have all passed the National Uniform Legal Profession Qualification Examination and are familiar with the Four-Element Theory. The entire annotation process took a total of 7 months and involved 4 rounds of annotation according to the validity of the legal source from high to low level.
The First Level: Article.
Legal elements can be seen as an interpretation and refinement of the statutory provisions corresponding to a particular crime. Using literal interpretation as the primary method, the statute is broken down based on its semantic meaning and common usage, ensuring that the interpretation does not extend beyond the possible meaning of the text: (1) linguistic analysis follows the subject-predicate-object structure of the provision. (2) To maintain consistency, terms are systematically classified and mapped(e.g: subjective aspect is classified as either intentional or negligent. )(3) Only when it is impossible to make an explicit inclusion or exclusion judgment for an element based on the rules of language use (neutral option field), other interpretation methods should be used. This initial phase takes almost 2 months.
For example, in the crime of robbery, the object "public or private property" represents the protected legal interest. The phrase "forcibly seizing public or private property through violence, coercion, or other means" describes the objective aspect. Since no subject is specified, it is assumed to involve a general subject, and the adverbs "violence" and "coercion" indicate an intentional act. Preliminarily interpret ‘violence’ in the objective aspect as ‘Use of physical force or power’, but the specific forms and subjects of violence need further clarification.
The Second Level: Judicial Interpretation.
In the 3rd and 4th months, the second level of the hierarchical annotation path focuses on refining legal elements through judicial interpretation. The primary method used for interpreting these materials is systematic interpretation. This approach examines the position of the corresponding articles within the legal system by analyzing their placement within the structure of laws, including parts, chapters, sections, articles, clauses, and subclauses, as well as their relationship to other statutes and judicial interpretations. Additionally, other interpretative methods, such as sociological interpretation and teleological interpretation, are referenced based on judicial interpretations, related statutory provisions, or bar exam questions.The goal of this level is to clarify the legislative intent by considering the contextual relevance of each provision within the broader legal framework.
For example, in the first level, the objective aspect of "violence" in the crime of robbery requires further clarification, specifically regarding whether violence must be directed exclusively at persons or could also apply to property. Article 289 of Chinese Criminal LawCongress (2017) stipulates that in cases of "smashing, looting, and robbing" committed by a group, the ringleaders shall be convicted of robbery if they destroy or seize public or private property. This provision demonstrates that violence against property can also constitute robbery under Chinese law.
The Third Level: Guiding Cases.
In the 5th to 6th month, purposive interpretation and sociological interpretation are applied to the guiding cases and landmark judgments from the Supreme Court. By examining the social significance of real-world cases, these methods bridge the subtle gap between abstract legal theory and practical cases. This approach enables dynamic adaptation and integration of empirical insights and emerging controversies within the dataset.
For example, in Criminal Trial Reference Case No.159Zou (2002), the perpetrator lured the victim into a room, locked the door, and seized 170,000 RMB intended for a transaction. The court determined that although the detention did not endanger personal safety, it was sufficient to suppress the victim’s resistance, thus constituting "violence" in in the objective aspects of robbery. Another example is the "Molestation and Theft Case" Ma (2021), where the perpetrator bound the victim, committed molestation, and stole the victim’s phone. Since the ongoing molestation reinforced coercion, it constitutes a new act of violence. Thus, the annotation includes "molestation" as an additional method.
The Fourth Level: Academic Discourses.
In the 7th month, the final stage involves academic expansion. Academic controversies are introduced by employing multiple interpretive methods such as comparative law interpretation, purposive interpretation, and sociological interpretation. These methods include inserting conflict markers at key points of controversy, highlighting the distinctions between mainstream consensus and minority theories, while providing brief annotations of their legal reasoning. This approach ensures the extensibility and academic depth of the dataset.
For example, regarding the crime of robbery, for the main view in China, Soviet Union, North Korea, and Japan explicitly holds that the violence must be severe enough to endanger the victim’s life or healthZhang (2007). But some scholars argue that any violence that can forcibly impact the victim’s body is sufficient to constitute violence in robbery, no need to endanger the victim’s life or healthYang (2010).
4 Data Distribution
Metric LLM Expert
Mean Median Mean Median
Avg. Length 115.43 - 472.53 -
SB 23.12 27 51.64 17
OB 15.86 15 36.01 25
SA 28.00 30 42.38 21
OA 48.45 45 342.5 230
Table 1: Comparison of Legal Element Lengths: LLM vs. Expert. SB = Subject, OB = Object, SA = Subjective Aspect, OA = Objective Aspect.
As shown in Table 1, we compare the length of legal elements between expert-annotated descriptions in JUREX-4E and LLM-generated outputs across 105 charges that overlap with the Lecard-V2 dataset Li et al. (2024c), which is one of the most comprehensive legal datasets, covering 184 criminal charges. We find that:
(1) The average total length of expert annotations (472.53) is more than four times longer than that of LLM-generated outputs (115.43), indicating that the former include more detailed information.
(2) The median difference between the Subject (SB), Object (OB), and Subjective Aspect (SA) is relatively small, as these elements are typically fixed. For example, the SB is often a general entity, and the SA is often intent or negligence.
(3) The median and mean values for SB and SA in the expert annotations differ, especially for SB (17 v.s. 51.64). This discrepancy arises because certain specialized charges may require more detailed explanations. For example, in the crime of copyright infringement, the definition of “work” under the subject element has 9 occasions. Detailed data distribution for each element is provided in Appendix A.
(4) The main difference between Expert and LLM is in the Objective Aspect (342.5 v.s. 48.45 in Mean). This is because the OA includes a range of factual elements describing the criminal behavior, such as the conduct, object, result, time, and location, which are most emphasized in legal provisions and are central to various legal interpretive theories.
5 Human Evaluation
We selected 6 complicated crimes in Chinese judicial practiceOuyang et al. (1999) to evaluate whether the LLM can handle the Four-element Theory. Drawing from previous workDeng et al. (2023); Cui et al. (2024); Zhou et al. (2023), we define LLM-generated knowledge as information produced by the LLM based on its pre-trained knowledge and contextual prompts. For detail, we provide the LLM with legal articles and the definition of each element in FET, prompting it to generate the four-elements base on these metrical. The LLM is expected to autonomously identify and generate the four elements based on its learned understanding of legal concepts.
We invite legal experts to assess the four elements generated by the LLM from four dimensions: Precision, Completeness, Representativeness, and Standardization.
  • Precision: Whether the key components of each element are accurately identified. This dimension mainly evaluates whether the four elements faithfully represent the legal provisions.
  • Completeness: Whether all necessary information of each element is included. This assesses whether any essential content is missing, such as the omission of a description for specific subjects, like government officials.
  • Representativeness: Whether the annotations highlight the most critical scenarios in judicial practice. For example, in crimes of intentional injury, this would involve describing the representative means of harm.
  • Standardization: Whether the four elements are clearly defined, ensuring consistency in the expression of identical elements across different crimes (e.g., consistent description of general subjects), with concise and easily understandable explanations, free from legal ambiguities or misunderstandings.
Each dimension was scored by two types of experts: one group with a pure legal background and another group with a combined background in law and Artificial Intelligence, all of whom have passed the bar examination. The experts were selected to balance domain expertise and interdisciplinary perspectives. Scores were averaged across the two groups. Details about 1-5 scale criteria and annotator background are provided in Appendix C.
Dimension LLM Expert $\delta$
Precision 4.12 4.69 + 0.57
Completeness 3.79 4.65 + 0.86
Representativeness 3.60 4.48 + 0.88
Standardization 4.33 4.56 + 0.23
Table 2: Performance comparison of four elements across methods. $\delta$ represents the score difference between expert and LLM-generated four-elements, with experts outperforming LLMs in all dimensions.
As shown in Table 2, expert annotations consistently outperform LLM-generated elements across all four dimensions, highlighting the limitations of LLMs in understanding legal elements. The most pronounced deficiencies are observed in Completeness (+0.86) and Representativeness (+0.88). This suggests that while LLMs can generate formally standardized and relatively accurate four elements, their description are not specific enough and do not adequately reflect the representative features of a charge’s criminal composition.
6 Evaluate Expert Knowledge on Charge Disambiguation
In the preceding section, the human evaluation demonstrated that experts annotated higher-quality four-elements. To further quantitatively evaluate the annotations, a direct way is to judge whether different charges can be distinguished according to the four-element definition of crime constitution. Therefore, we introduce the Similar Charge Disambiguation (SCD) taskYuan et al. (2024); Li et al. (2024a).
6.1 Experiment Settings
6.1.1 Dataset
We chose the dataset released by Liu et al. (2021), which includes five charge sets with the largest number of cases. To evaluate performance on representative tasks, we selected three 2-label classification groups commonly examined in other datasets Yuan et al. (2024): Fraud & Extortion (F&E), Embezzlement & Misappropriation of Public Funds (E&MPF), and Abuse of Power & Dereliction of Duty (AP&DD). Each crime has over 1.9k cases, with a total of 13,962 cases. The details of the classification groups are shown in Appendix D. Following previous work Liu et al. (2021); Yuan et al. (2024), we use Average Accuracy (Acc) and macro-F1 (F1) as evaluation metrics.
6.1.2 Baselines and Methods
To evaluate SCD tasks, we consider two ways of incorporating legal knowledge. The first directly integrates legal statutes, represented by GPT-4o Achiam et al. (2023) as the baseline and GPT-4o+Article, which explicitly provides relevant legal articles to the model. The second adopts structured legal reasoning to enhance interpretability and accuracy. We consider Legal-CoT, a Chain-of-Thought Kojima et al. (2022) variant that conducts a stepwise analysis based on the FET, and MALR Yuan et al. (2024), a multi-agent framework that decomposes legal tasks into sub-tasks in four-element structures. Details of each baseline are provided in Appendix 7.
We use an unified approach to introduce four-element descriptions. For each group of similar charges, the model receives charges’ four-elements from JUREX-4E or generated by LLM to aid classification. Specifically, GPT-4o+FETExpert relies on expert-annotated four-elements, while GPT-4o+FETLLM relies on LLM-generated four-elements. As shown in Appendix D, the instruction format is consistent across methods, with only the [Four Elements of candidate charges] varying based on the source. All experiments are conducted in a zero-shot setting, with the max_tokens set to 3,000 (or 10,000 for COT and MALR reasoning) and temperature set to 0 or 0.0001(In repeated experiments).
6.2 Results
Model F&E E&MPF AP&DD Average
Acc F1 Acc F1 Acc F1 Acc F1
GPT-4o 94.36 95.81 86.49 89.76 85.54 87.12 88.72 90.07
GPT-4o+Article 95.34 96.30 92.64 93.03 88.30 89.33 92.09 92.89
Legal-COT 94.99 96.27 90.50 90.99 87.81 88.14 89.95 90.85
MALR 94.62 95.82 86.99 86.98 87.86 88.68 89.82 90.49
GPT-4o+FETLLM 95.73 96.56 91.87 92.01 89.61 89.69 92.40 92.75
GPT-4o+FETExpert 96.06 96.69 92.57 93.05 90.53 90.62 93.05 93.45
Table 3: Results of Charge Disambiguation. FET means introducing the Four-element theory with knowledge obtained from experts and LLM method. Highest results are in bold.
As shown in Table 3, the GPT-4o+FETExpert performs best in discriminating similar charges, indicating that expert annotation is superior to other methods of directly or indirectly introducing FET with LLMs. Specifically, we can derive the following observation:
Effectiveness of Domain-Specific Legal Knowledge:
Among all approaches, those that explicitly incorporate domain-specific legal knowledge, such as GPT-4o+Article, Legal-CoT, and MALR, outperform GPT-4o alone. This highlights the importance of integrating legal knowledge.
Importance of Concrete Four-element Knowledge:
The accuracy of both Legal-CoT and MALR is still lower than GPT-4o+FET methods. This suggests that, compared to embedding the Four-Element Theory into LLMs’ reasoning process, providing concrete charge four-elements enables the model to better understand the different crimes’ composition.
Superiority of Expert Annotations:
Compared with the indirect introduction of FET reasoning, the method of directly introducing four-elements to the model (GPT-4o+FET) achieves better results. Notably, GPT-4o+FETExpert surpassing the GPT-4o+FETLLM by 0.65 in average accuracy and 0.70 in average F1-score, underscoring the superior quality and reliability of expert annotations in legal tasks, aligning with human evaluations in Table 2 and reaffirming the critical role of human expertise in legal decision-making.
7 Can Expert Knowledge Benefit More Downstream Tasks?
In this section, we design a simple framework to apply the expert-annotated four elements to Legal Case Retrieval (LCR), a task in which relevant cases are retrieved based on given facts. It is an important step in the practice of analyzing cases and making judgments, and it requires the precise application of the four-element theory to matches cases with similar criminal compositions.
Figure 2: An expert-guided FET method to enhance legal case retrieval by incorporating expert four elements.
7.1 Method
Model NDCG@10 NDCG@20 NDCG@30 R@1 R@5 R@10 R@20 R@30 MRR
BERT 0.1511 0.1794 0.1978 0.0199 0.0753 0.1299 0.2157 0.2579 0.1136
Legal-BERT 0.1300 0.1487 0.1649 0.0186 0.0542 0.1309 0.1822 0.2172 0.0573
Lawformer 0.2684 0.3049 0.3560 0.0432 0.1479 0.2330 0.3349 0.4683 0.1096
ChatLaw 0.2049 0.2328 0.2745 0.0353 0.1306 0.1913 0.2684 0.3751 0.1285
SAILER 0.3142 0.4133 0.4745 0.0539 0.1780 0.3442 0.5688 0.7092 0.1427
GEAR * * * 0.0630 0.1706 0.3142 0.4625 * 0.2162
BGE 0.4737 0.5539 0.5937 0.0793 0.2945 0.4298 0.6500 0.7394 0.1926
FETLLM 0.5139 0.5862 0.6291 0.0980 0.2967 0.4769 0.6802 0.7828 0.2140
        - base 0.3583 0.4293 0.4798 0.0506 0.2240 0.3644 0.5383 0.6652 0.1453
FETExpert_guided 0.5211 0.5920 0.6379 0.1024 0.3049 0.4883 0.6885 0.7967 0.2155
        - base 0.3766 0.4584 0.5111 0.0715 0.1894 0.3709 0.5891 0.7203 0.1624
Table 4: SCR results. Bold fonts indicate leading results in each setting. * denotes that the indicator is not applicable to the current model.
We implement a standard dense retrieval approach BGE using BGE-m3 Chen et al. (2023), an advanced embedding model for dense retrieval. Given a query $q$ and a candidate case $c$, their vector representations $\mathbf{v}_{q}$ and $\mathbf{v}_{c}$ are obtained through shared encoder $E$: $\mathbf{v}_{q}=E(q),\quad\mathbf{v}_{c}=E(c)$. We used the BGE-m3 model without fine-tuning as the shared encoder. Next, the relevance score is computed via cosine similarity:
$\text{sim}_{\text{base}}(q,c)=\frac{\mathbf{v}_{q}\cdot\mathbf{v}_{c}}{\|\mathbf{v}_{q}\|\|\mathbf{v}_{c}\|}$ (1)
To retrieve the top-k most similar cases, we rank the candidates based on their cosine similarity to the query. Denote the set of candidate cases as $C=\{c_{1},c_{2},\dots,c_{n}\}$, where $n$ is the total number of candidate cases. We compute the similarity for each $c_{i}\in C$, and select the top-k candidates with the highest similarity scores.
As shown in Figure 2, to leverages expert-annotated four elements of charges, we introduce an BGE+FETExpert_guided method for the retrieval process, consisting of three steps: (1) Predicting charges, a LLM $\mathcal{M}_{p}$ predicts potential charges $Z=\{z_{1},...,z_{k}\}$ from case facts. (2) Matching elements, retrieving corresponding charge’s four-elements $\{f_{z}\}_{z\in Z}$ in JUREX-4E. (3) Analyzing case facts. Guided by $\{f_{z}\}$, another LLM $\mathcal{M}_{g}$ generates case-specific four elements $a_{c}$ for candidate $c$. The final similarity score combines factual and theoretical alignment:
$\text{sim}_{\text{final}}(q,c)=\alpha\cdot\text{sim}_{\text{base}}(q,c)+(1-\alpha)\cdot\text{sim}_{\text{f}}(a_{q},a_{c})$ (2)
where $\alpha=0.7$ and $\text{sim}_{\text{f}}$ measures the similarity between the generated four-element descriptions.
To facilitate comparison, we also design a BGE+FETLLM method that directly prompt the LLM $\mathcal{M}_{g}$ with the concept of Four-Element Theory to generate case-specific four elements $a_{c}$.
7.2 Dataset
LeCaRDv2Li et al. (2024c) is the latest version of LeCaRDMa et al. (2021), which is widely used in LCR task Li et al. (2024b); Zhou et al. (2023). It comprises 800 queries and 55,192 candidates extracted from 4.3 million criminal case documents. There are two common evaluation settings for this dataset: one uses a subset Qin et al. (2024) with a candidate pool size of 1,390, while the other uses the full set Li et al. (2024c) with a candidate pool size of 55,000. We conducted experiments under both settings.
Following previous workFeng et al. (2024); Qin et al. (2024), we adopt commonly used evaluation metrics. For the subset, we use NDCG@10, 20, 30, Recall@1, 5, 10, 20, and MRR. For the full dataset, we use Recall@100, Recall@200, Recall@500, and Recall@1000.
7.3 Baselines
Consistent with earlier workLi et al. (2024c); Qin et al. (2024), we compare some dense retrieval methods, including: BERTDevlin (2018), LawformerXiao et al. (2021), ChatLaw-Text2Vec [^1] https://modelscope.cn/models/fengshan/ChatLaw-Text2VecCui et al. (2023), SAILERLi et al. (2023), GEARQin et al. (2024). Details of each baseline is shown in Appendix E. These baselines are implemented using the FlagEmbedding Toolkit [^2] https://github.com/FlagOpen/FlagEmbedding with a RTX 3090.
7.4 Results
The LCR results are shown in Table 4, where we can observe that:
FET Works Well in LCR.
The baseline model BGE achieves strong performance across most metrics compared to previous methods. Introducing the Four-Element Theory (FET) further improves its results, with relative MRR improvements of 11.11% for FETLLM and 11.89% for FETExpert_guided, indicating that introducing legal theory is effective.
Expert Knowledge is Necessary.
By leveraging external knowledge, FETExpert_guided achieves significant improvements across all of the metrics. Specifically, using expert-guided case four-elements (FETExpert_guided-base) outperforms LLM-generated case four-elements (FETLLM-base) by an average of 11.77% in MRR, demonstrating the critical role of expert knowledge in enhancing retrieval precision. A case study in Appendix G shows that the expert four-element for charges provide practical judgment points and key narratives (e.g., the special subject of the Crime of Embezzlement) that help the LLM focus on essential facts to analyze the case.
We also evaluated the FET method on the full set, as shown in Table 9 , and the results remain consistent, with the expert-guided method still performing best.
8 Conclusion
In this paper, we propose an expert-annotated knowledge base, evaluate its quality in the Similar Charge Distinction task, and apply it to the Legal Case Retrieval task. Our results demonstrate that expert annotations significantly enhance LLMs’ understanding of the Four-Element Theory. The four-element annotations, enriched with professional legal interpretations, provide strong support for LLMs’ reasoning capabilities. This approach can be extended to other legal AI tasks, such as legal document analysis and contract interpretation.
9 Ethical Considerations
The datasets used in our evaluation are sourced from publicly available legal datasets, with all defendant information anonymized to ensure privacy.
10 Limitations
As a limitation, this knowledge base focuses on the Four-Element Theory within the context of 155 crimes under Chinese Criminal Law. However, the four-level hierarchical pyramid annotation structure based on the legal interpretation system proposed in this work provides valuable insights for future expansion to other legal domains, as it represents a theoretical framework in the field of jurisprudence. The interpretative methods within the legal interpretation system, including textual, systematic, sociological, and doctrinal interpretations, are universally recognized in international law field and can be applied to different laws, countries, and legal systems.
Acknowledgments
Appendix A Detailed Data Distribution for each Element
Figure 3: The average length distribution of total four elements annotated by experts.
Figure 4: The length distribution of each element annotated by experts.
Figure 5: The average length distribution of total four elements generated by LLM.
Figure 6: The length distribution of each element generated by LLM.
Appendix B Interpretation Methods
1. Literal Interpretation
A strict textual analysis method that adheres to the ordinary meaning of words as understood by a reasonable person at the time of enactment, excluding subjective intent inference
2. Systematic Interpretation
An approach interpreting legal provisions through their position within the codified legal hierarchy and logical connections with related norms, maintaining the integrity of the legal system (aligned with Dworkin’s "law as integrity" theory).
3. Purposive Interpretation
A method discerning the objective legislative purpose through analysis of statutory structure and functional goals, distinct from subjective legislative intent (following Hart & Sacks’ legal process school).
4. Historical Interpretation
Interpretation based on legislative history materials including drafts, debates and official commentaries, while distinguishing original meaning from framers’ subjective intentions (as per Brest’s original understanding theory).
5. Comparative Interpretation
A methodology referencing functionally comparable legal systems sharing common juridical traditions, employing analogical reasoning while considering local legal culture (developed through Gottfried Wilhelm Leibniz’s comparative law framework).
6. Sociological Interpretation
Interpretation evaluating social efficacy through empirical analysis of implementation effects, guided by Pound’s sociological jurisprudence principle that "law must be measured by its achieved results".
Appendix C Human Evaluation Guidance
The annotators included three postgraduate students specializing in criminal law and one master’s student in legal science and technology. The annotators scored independently, without knowledge of each other’s results. Before scoring, they were asked to read the descriptions and scoring guidelines (as shown in Table 5) for each evaluation dimension. In order to ensure the fairness of the evaluation, they do not know the source of each four elements, and even do not know that these four elements include those generated by LLMs.
When assigning scores, they were also required to provide brief justifications. For example, for the Completeness dimension: 3 (The description of Objective Aspect is too brief, and does not specify the intent of illegal possession).
Dimension Precision Completeness Representativeness Standardization
Definition Whether there are errors in key elements Whether the four elements are complete Whether key elements and scenarios are emphasized Whether language and format are clear and standardized
Score 1 Contains numerous obvious errors, severely impeding the judgment of culpability, exculpation, and conviction, leading to significant deviations. Severe omission of key content, unable to present a complete picture of the crime structure, greatly hindering analysis of criminal behavior. Completely fails to mention any key elements or scenarios, unable to highlight essential points for crime recognition, offering no assistance in conviction. Language is extremely chaotic and obscure; format lacks any standardization, greatly hindering comprehension and application.
Score 2 Contains multiple noticeable errors, significantly interfering with culpability, exculpation, and conviction judgments, potentially leading to partial errors. Noticeable omissions in content, failing to comprehensively cover crime elements, affecting thorough analysis of criminal behavior. Only highlights a minimal and unimportant portion of the key elements, providing weak support for understanding key crime features. Language is relatively vague and inaccurate, with a casual format that makes content comprehension significantly challenging.
Score 3 Contains a few errors, but the overall accuracy in determining culpability, exculpation, and conviction is relatively unaffected, unlikely to lead to judgment errors. Some key content descriptions are incomplete, but they generally present the framework of the crime structure. Highlights some relatively important key elements but lacks comprehensiveness and prominence, offering limited assistance in crime identification. Language is generally clear but may have minor deviations in phrasing or formatting.
Score 4 Almost error-free, key elements accurately serve culpability, exculpation, and conviction judgments, ensuring the accuracy of results. Key elements are mostly complete, with only very slight and non-critical deficiencies that do not hinder a comprehensive analysis of the crime. Clearly and relatively comprehensively highlights key elements, aiding in accurately identifying crucial aspects of criminal behavior. Language is clear and accurate, format is relatively standardized, facilitating comprehension and application of relevant content.
Score 5 Completely error-free, key elements are precisely defined, achieving highly accurate culpability, exculpation, and conviction judgments without any flaws. All four elements are complete and detailed, covering every aspect of the crime, perfectly presenting the crime structure. Precisely and comprehensively highlights all crucial elements, enabling immediate grasp of the core aspects of the crime, significantly aiding conviction. Language is extremely clear, standardized, and concise; format perfectly meets requirements, with no barriers to understanding, ensuring efficient information delivery.
Table 5: The four dimensions of the human evaluation and the specific score description.
Appendix D Details for Similar Charge Disambiguation
Charge Sets Charges Cases
F&E Fraud & Extortion 3536 / 2149
E&MPF Embezzlement & Misappropriation of Public Funds 2391 / 1998
AP&DD Abuse of Power & Dereliction of Duty 1950 / 1938
Table 6: Distribution of charges in the GCI dataset. Cases denotes the number of cases in each category. Following Liu et al. (2021), for a case with both confusable charges, the prediction of any one of the charges is considered correct.
Prompt:
You are a lawyer specializing in criminal law. Based on Chinese criminal law,
please determine which of the following candidate charges the given facts align with.
The candidate charges and their corresponding four elements are as follows:
[Four Elements of Candidate Charges].
The four elements represent the core factors for determining the constitution of a criminal charge.
[The basic concepts of the Four-Element Theory]
Please Compare the case facts to determine which charge’s four elements they align with, thereby identifying the charge.
Table 7: Prompt template for adding the Four-Element Theory and specific four elements of crime in charge disambiguation.
For LLM baselines, we evaluate both general-purpose and task-specific methods.
GPT-4o is an optimized version of GPT-4Achiam et al. (2023) that has well performance in specific tasks through domain adaptation.
To explore the effectiveness of notes-guided four elements in LLMs, we further consider other methods that introduced the Four-element theory into LLMs.
GPT-4oLaw, which introduces articles related to corresponding charges into the instruction to provide legal context.
Legal-COT is a variant of COT Kojima et al. (2022) that guides the LLM to perform step-by-step legal reasoning by incorporating explanations of the Four-element theory into the instruction.
MALR is a up to date multi-agent framework designed to enhance complex legal reasoning Yuan et al. (2024), enabling LLMs to autonomously decompose legal tasks and extract insights from legal rules. As its full implementation is not publicly available, we use the released code for the auto-planner module and implement the legal insight extraction following the specified steps and prompts, with necessary refinements. Experiments on the paper’s reported examples show that our implementation produces task decompositions and outputs largely consistent with the original results.
As shown in Table D, different methods differ in their prompts for generating and explaining the Four-Element Theory, but generally follow a similar process. For the SCD output, except for COT and MALR, which require reasoning processes and prediction results, all other methods only require the output of prediction results.
Method GPT-4o GPT-4o+Article Legal-COT GPT-4o+FETLLM GPT-4o+FETExperts
Pre-task None None None LLM-generated four elements Expert-annotated four elements
Prompt You are a lawyer specializing in criminal law. Based on Chinese criminal law, please determine which of the following candidate charges the given facts align with.
Candidate charges are as follows: #Candidate Charges The candidate charges and relevant legal articles are as follows: #Candidate Charges + #Articles Please analyze using the Four Elements Theory step by step: #details about each step. The candidate charges are as follows: #Candidate Charges The candidate charges and their corresponding four elements are as follows: #Four Elements of candidate charges. The four elements represent the four core factors of a charge. Compare the case facts to determine which charge’s four elements they align with, thereby identifying the charge.
Output format: #Format. Note: Only output the charge, no additional information.
Case facts: #Case Facts.
Table 8: Prompts of different methods in Similar Charge Disambiguation. # represents a format input.
Appendix E Baselines in Legal Case Retrieval
BERTDevlin (2018) is a language model widely used in retrieval tasks. In this paper, we chose BERT-base-Chinese [^3] https://huggingface.co/google-bert/bert-base-chinese. Legal-BERT [^4] https://github.com/thunlp/OpenCLaPChalkidis et al. (2020) is a variant of BERT that is specifically trained on legal corpora. LawformerXiao et al. (2021)is a Chinese legal pre-trained model based on LongformerBeltagy et al. (2020), which is able to process long texts in the legal domain. ChatLaw-Text2Vec [^5] https://modelscope.cn/models/fengshan/ChatLaw-Text2VecCui et al. (2023) is a Chinese legal LLM trained on 936,727 legal cases for similarity calculation of legal-related texts. SAILERLi et al. (2023) is a structure-aware legal case retrieval model utilizing the structural information in legal case documents. GEARQin et al. (2024) is a generative retrieval framework that explicitly integrates judgment prediction with legal document retrieval in a sequence-to-sequence manner. Since the output of GEAR cannot directly evaluate NDCG, the official results under the same setting are directly referenced in this paper. LLM and Expert represent the results of retrieval using only the four elements.
Appendix F SCR results on the full LeCaRDv2 Dataset
As presented in Table 9, we selected several representative methods based on sparse retrieval and dense retrieval for experiments on the full LeCaRDv2 dataset. All language models were not fine-tuned. The notes-guided FET method achieved the best performance among all language models, attaining top results in both R@500 and R@1000. The results indicate that the conclusions drawn from the full dataset are consistent with those from the subset, and the notes-guided method demonstrates strong performance.
Model R@100 R@200 R@500 R@1000
BERT 0.1116 0.1493 0.2174 0.2819
Lawformer 0.2432 0.304 0.4054 0.4833
ChatLaw 0.1045 0.1628 0.2791 0.3999
SAILER 0.2834 0.4033 0.6104 0.7568
BGE 0.4085 0.5246 0.6855 0.7912
FETLLM 0.4167 0.5388 0.7006 0.7925
FETExpert_guided 0.4201 0.5396 0.7010 0.7927
Table 9: SCR results on the full set of LeCaRDv2. Bold fonts indicate leading results in each setting. The expert-guided FET method achieved the best performance among all language models and attained the top results in both R@500 and R@1000.
Appendix G A Case Study of LCR
Table 10 presents a case study on the Crime of Embezzlement. By comparing the four elements annotated by experts for the crime in JUREX-4E, the case-specific four elements generated directly by the LLM, and those generated by the LLM with expert four elements of charge as guidance, we can observe that:
1) Incorporating expert fine-grained annotations enables the model to better grasp the elements of a crime, thereby providing more precise element comparison. For example, LLMs can identify the “integrity of official duties”, and the subjective aspect “Intentional” can be interpreted as “having the purpose of illegally possessing public or private property”, highlighting the characteristics of “official duties”. Capturing the core information of the case is crucial for matching cases with similar facts.
2) LLMs can conduct case-tailored specific analysis based on the constitutive elements of a crime. Blue parts show the LLMs can better analyze the defendant’s workplace and the actions taken in the case, which reflects the significance of specific and accurate legal knowledge.
Document [Head of document]…In April 201X, Company A appointed B as the Sales Manager and Deputy Manager of the Catering Department, responsible for collecting outstanding debts. From December 201X to April 201Y, B embezzled X yuan from six units’ payments for personal investment and consumption. In May 201Y, knowing that the company required him to reconcile accounts with the finance department, B fled after changing his phone number. It was found that after the case broke out, B’s father C had returned the X yuan to Company A.
Charge Crime of Embezzlement
Experts Four-elements LLM-generated Experts-guided
Four Elements
Object: The ownership of property and integrity of duties in companies, enterprises, or other organizations. The objects include movable and immovable, tangible and intangible property of the perpetrator’s organization.
Objective Aspect: Using positional convenience to illegally appropriate organizational property in a significant amount.
Subject: Special subject, namely personnel of companies, enterprises, or other organizations.
Subjective Aspect:Intentional, with the purpose of illegally possessing property.
Object: The ownership of property of companies, enterprises, or other organizations.
Objective Aspect: Using the convenience of one’s position to illegally appropriate the property of the organization, in a significant amount.
Subject: Personnel of companies, enterprises, or other organizations. The defendant is B.
Subjective Aspect: Intentional.
Object: The ownership of property of companies, enterprises, or other organizations and the integrity of official duties. In this case, crime objetcs are the properties of the company, enterprise, or other organization where the perpetrator is employed.
Objective Aspect: The act of using one’s position for personal gain by illegally appropriating the property of the organization, where the amount involved is significant. Specifically, B used his position to embezzle X yuan from the payments of six units for personal investment and consumption.
Subject: Special subject, namely personnel of companies, enterprises, or other organizations. Defendant B is the Sales Manager and Deputy Manager of the Catering Department of the company.
Subjective Aspect: Intentional, with the purpose of illegally possessing public or private property.
Table 10: The results of the case four elements obtained through different methods in LCR. Notes refer to expert notes related to the charge retrieved during the search. LLM-generated and Experts-guided indicate whether using JUREX-4E’s four-elements of the crime to guide LLM in generating the four elements. Red parts mean the knowledge from JUREX-4E, while blue parts show the LLM’s internal knowledge. By incorporating JUREX-4E, the model better emphasizes conviction and sentencing related information and provides more detailed descriptions of critical case facts.

### v2

@@
 Abstract
-The Four-Element Theory is a fundamental framework in criminal law, defining the constitution of crime through four dimensions: Subject, Object, Subjective aspect, and Objective aspect.
-This theory is widely referenced in legal reasoning, and many Large Language Models (LLMs) attempt to incorporate it when handling legal tasks.
-However, current approaches rely on LLMs’ internal knowledge to incorporate this theory, often lacking completeness and representativeness.
-To address this limitation, we introduce JUREX-4E, an expert-annotated knowledge base covering 155 criminal charges.
-It is structured through a progressive hierarchical annotation framework that prioritizes legal source validity and employs diverse legal interpretation methods to ensure comprehensiveness and authority.
-We evaluate JUREX-4E on the Similar Charge Distinction task and apply it to Legal Case Retrieval, demonstrating its effectiveness in improving LLM performance.
+In recent years, Large Language Models (LLMs) have been widely applied to legal tasks.
+To enhance their understanding of legal texts and improve reasoning accuracy, a promising approach is to incorporate legal theories.
+One of the most widely adopted theories is the Four-Element Theory (FET), which defines the crime constitution through four elements: Subject, Object, Subjective Aspect, and Objective Aspect.
+While recent work has explored prompting LLMs to follow FET, our evaluation demonstrates that LLM-generated four-elements are often incomplete and less representative, limiting their effectiveness in legal reasoning.
+To address these issues, we present JUREX-4E, an expert-annotated four-element knowledge base covering 155 criminal charges.
+The annotations follow a progressive hierarchical framework grounded in legal source validity and incorporate diverse interpretive methods to ensure precision and authority.
+We evaluate JUREX-4E on the Similar Charge Disambiguation task and apply it to Legal Case Retrieval.
 Experimental results validate the high quality of JUREX-4E and its substantial impact on downstream legal tasks, underscoring its potential for advancing legal AI applications.
-Code: https://github.com/THUlawtech/JUREX
+The dataset and code are available at: https://github.com/THUlawtech/JUREX
 JUREX-4E: Juridical Expert-Annotated Four-Element Knowledge Base for Legal Reasoning
@@
 1 Introduction
-In legal AI tasks, enhancing the accuracy and interpretability of Large Language Models (LLMs) in the legal domain often requires the incorporation of legal theories as a support Jiang and Yang (2023); Servantez et al.
+Large Language Models (LLMs) have recently demonstrated impressive performance in legal tasks such as charge prediction Yuan et al.
+(2024) and legal case retrieval Feng et al.
+(2024).
+In these applications, a key challenge is accurately understanding complex legal language.
+To address this, recent studies have introduced legal theories into LLM workflows Jiang and Yang (2023); Servantez et al.
 (2024); Yuan et al.
 (2024); Deng et al.
+(2023), as these theories provide structured reasoning frameworks and domain knowledge.
+Among these theories, the Four-Element Theory (FET) in Chinese criminal law Liang (2017) is particularly important, as it defines the legal criteria for establishing criminal liability.
+FET breaks down a criminal charge into four elements: Subject, Object, Subjective Aspect, and Objective Aspect, which serve as the essential criteria for determining whether a defendant’s behavior constitutes a specific crime.
+Most current approaches rely on the LLM’s internal knowledge to incorporate the FET.
+A common method is to ask LLMs to emulate expert reasoning processes.
+For example, designing four separate prompts to guide the LLM outputs in the form of four-elements Deng et al.
 (2023).
-One important theory is the Four-Element Theory of Crime Constitution in Chinese criminal law Liang (2017).
-This theory deconstructs criminal conduct into four elements: Subject, Object, Subjective aspect, and Objective aspect, providing clear standards for judicial authorities to determine criminal behavior and helping to prevent the abuse of penal power.
-However, most current approaches do not provide additional knowledge but rather rely on the LLM’s internal knowledge to incorporate the Four-Element Theory.
-A common method is to guide LLMs in mimicking expert reasoning processes.
-For example, designing four separate prompts to guide the LLM outputs in the form of four elementsDeng et al.
+This raises a critical question: Can LLMs reliably understand and apply the FET?
+Figure 1: An example of LLM-generated four-elements.
+To investigate this, we conducted a pilot study where we provide LLMs with legal articles and asked them to generate the four-elements for several representative charges Ouyang et al.
 (1999).
-These methods assume that the model has a solid grasp of the Four-Element Theory, which has not yet been verified.
-We had LLMs generate the four elements of several complicated crimes in Chinese judicial practiceOuyang et al.
-(1999), and then asked legal experts to score them.
-We found that, although LLMs can generate formally standardized and relatively accurate legal descriptions when provided with legal theoretical frameworks and references, the model still underperformed in terms of completeness and representativeness.
-This shortcoming could affect the accuracy and soundness of subsequent reasoning.
-To help LLMs better utilize the Four-Element Theory in legal tasks, we propose JUREX-4E: JURidical EXpert-annotated 4-Element knowledge base for legal reasoning.
-This knowledge base is annotated using a progressive hierarchy: Article → Judicial Interpretations → Guiding Cases → Academic Discourses, which is built upon the pyramid structure of legal source validity.
-It incorporates various legal interpretation methods, including textual, systematic, sociological, and purposive interpretations.
-The knowledge base covers the four elements of 155 high frequency charges, annotated by legal experts over a period of seven months.
-Each crime’s four elements are described in an average of 472.5 words.
-To assess the quality of the annotations, we sampled several crimes for human evaluation.
-The expert annotations achieved an average score of 4.60 on a 5-point scale, while the LLM-generated four elements scored only 3.96, indicating that the expert annotations were of higher quality.
-To further evaluate the annotations objectively and comprehensively, a direct way is to judge whether different charges can be distinguished according to the four-element definition of crime constitution.
-Therefore, we introduced the Similar Charge Distinction task Liu et al.
-(2021).
-For each case, we provided the four elements of the candidate confused charges and combined them with the case facts as model input.
-The experimental results showed that injecting expert annotations helped the model better differentiate between similar charges, improving performance with a 0.65 increase in average accuracy and a 0.70 increase in average F1-score, underscoring the superior quality and reliability of expert annotations compared to those generated by the LLM.
-We also applied the expert annotations in a specific legal task: Legal Case Retrieval.
-It is an important step in the practice of analyzing cases and making judgments, requiring the precise application of the four-element theory to compare the criminal composition of cases.
-We designed a simple retrieval framework guided by expert knowledge, in which the charge’s four elements was used to generate four-element descriptions for both the query and candidate cases, and then match similar cases based on their vector similarity.
-Experiments demonstrated that incorporating expert-annotated four elements improved retrieval performance, as the model became better at focusing on the legal features and key details.
+Results show that the LLM-generated four-elements are often not accurate enough.
+As shown in Figure 1, in the charge of misappropriation of public funds, the LLM failed to identify the right to benefit from the use of public funds, a core part of the Object.
+These results suggest that LLMs lack the domain knowledge and legal reasoning precision required for reliable FET application.
+To help LLMs better utilize the FET in legal tasks, we construct JUREX-4E: JURidical EXpert-annotated 4-Element knowledge base for legal reasoning.
+The knowledge base covers 155 high-frequency criminal charges, each decomposed into Subject, Object, Subjective Aspect, and Objective Aspect.
+JUREX-4E is built through a four-stage Hierarchical Legal Interpretation System.
+In this process, legal experts refine each element by referencing sources in descending order of legal validity—Criminal Articles, Judicial Interpretations, Guiding Cases, and Academic Discourses—while applying appropriate interpretive methods at each stage.
+Each charge was annotated over a seven-month period, yielding knowledge-rich representations with an average annotation length of 472.5 words.
+To assess the quality of JUREX-4E, we conduct a human evaluation on four independent dimensions, Precision, Completeness, Representativeness, and Standardization, grounded in legal scholarship on how criminal elements should be normatively defined and expressed in judicial contexts  Zhang (2007a).
+The expert-annotated four-elements achieved an average score of 4.60 on a 5-point scale, significantly outperforming the LLM-generated ones, which scored 3.96.
+Among the four dimensions, the largest performance gaps appeared in Completeness and Representativeness, as expert annotations provided more comprehensive legal interpretations and summarized typical application scenarios, which are often overlooked by LLMs.
+To further evaluate the quality and utility of JUREX-4E, we conducted two downstream tasks: Similar Charge Disambiguation (SCD) and Legal Case Retrieval (LCR).
+In the SCD task Liu et al.
+(2021), we tested whether different charges could be more effectively distinguished by incorporating four-element knowledge.
+Results show that expert-annotated four-elements from JUREX-4E consistently outperformed LLM-generated counterparts across various prompting strategies and model types, improving average accuracy by 0.70% and F1-score by 0.75%.
+In the LCR task Li et al.
+(2024d), we incorporated JUREX-4E into the retrieval pipeline to guide case-level four-element generation and similarity matching, achieving better retrieval accuracy.
+Together, these findings validate the high quality and practical value of JUREX-4E in enhancing legal understanding and decision-making.
 Our contributions are as follows:
   1. (1)
-    We verify that LLMs have gaps in understanding the legal theory, highlighting the inadequacy of relying solely on LLM-driven reasoning for legal AI tasks.
+    We demonstrate that while LLMs can assist legal reasoning to some extent, they still fall short in accurately understanding and applying the Four-Element Theory.
   2. (2)
-    We built the JUREX-4E knowledge base, which is the first to incorporate the pyramid structure of legal source validity and covers the four elements of 155 criminal charges under Chinese Criminal Law.
+    We construct the JUREX-4E, the first expert-annotated legal knowledge base grounded in a hierarchical legal interpretation framework based on legal source validity.
   3. (3)
-    We demonstrated the significance of incorporating criminal composition elements in the Similar Charge Distinction task and proved the superior quality of the expert-annotated four-element knowledge base.
-  4. (4)
-    We applied JUREX-4E to the Legal Case Retrieval task, found that they do indeed contribute to downstream tasks.
+    We validate the quality and effectiveness of JUREX-4E on two representative legal tasks, Similar Charge Disambiguation (SCD) and Legal Case Retrieval (LCR), where it consistently outperforms LLM-generated representations across various prompting strategies.
+2 Background
+The Four-Element Theory (FET) of crime constitution is a fundamental framework in Chinese criminal law Liang (2017).
+It provides a standardized structure to determine criminal liability through four elements: Subject, Object, Subjective Aspect, and Objective Aspect.
+For example, the four-elements of Robbery can be briefly summarized as follows:
+(1) Subject (the person who commits a criminal act and should bear criminal responsibility): General subjects above the age of criminal responsibility.
+(2) Object (the legal interest harmed by the act): A compound object, combining both property ownership and personal rights of the victim.
+(3) Subjective Aspect (the offender’s mental state regarding the harmful act): Direct intent with the purpose of unlawfully appropriating another’s property.
+(4) Objective Aspect (the external facts of the criminal activity, including key actions and their outcomes): On-the-spot taking of property from an owner, custodian, or possessor through violence, coercion, or other methods.
+For the legal community, FET plays a central role in doctrinal analysis and judicial reasoning.
+It serves as the legal basis for both legislation and adjudication, ensuring internal consistency and normative rigor in criminal law application Li (2006); Zhang (2007a).
+For the legal AI community, FET offers a task-agnostic and interpretable framework for modeling legal reasoning Deng et al.
+(2023); Yuan et al.
+(2024).
+Compared to general reasoning templates (e.g., legal syllogism Gold (2018)) or alternative theories such as the Three-Tier System Zhou (2017); Zhang (2010), FET has become the dominant approach in China for assessing criminal liability Wang (2017).
+Its clearer and more interpretable decomposition of crimes into objective and subjective elements makes it particularly suitable for structured legal reasoning tasks.
 3 Related Work
-In legal AI, much work has introduced legal theories to enhance reasoning and improve model accuracy and interpretability.
-For example, legal syllogism prompting (LoT)Jiang and Yang (2023) teaches LLMs for legal judgment prediction by instructing legal syllogism, Chain of LogicServantez et al.
-(2024) guides models in reasoning about compositional rules by decomposing logical statements based on the IRAC (Issue, Rule, Application, Conclusion) paradigm.
-Among these, the Four-Elements Theory (FET) of Crime Constitution is a widely adopted frameworkYuan et al.
-(2024); Deng et al.
+With the rise of open-source base LLMs, lots of legal LLMs have emerged, such as Lawyer LLaMA Huang et al.
+(2023), DiSC-LawLLM Yue et al.
+(2023), ChatLaw Cui et al.
+(2024), and TongyiFarui [^1] https://tongyi.aliyun.com/farui.
+These models are typically adapted from general-purpose LLMs via domain-specific post-training or Retrieval-Augmented Generation (RAG), incorporating legal texts like cases and laws.
+Although these models achieve notable improvements on legal tasks, they still struggle with complex legal reasoning, such as charge disambiguation, legal question answering, statutory interpretation, and structured explanation generation Hu et al.
 (2025).
-The Four-Element Theory is one of the most widely recognized criminal theories in Chinese judicial practice Liang (2017).
-It specifies four essential elements that must be satisfied to establish criminal liability: Subject, Object, Subjective aspect, and Objective aspect.
-For example, the four elements of the Crime of Affray can be briefly summarized as follows:
-(1) Subject: Principal organizers and other active participants who have reached the age of criminal responsibility.
-(2) Object: Public order.
-(3) Objective Aspect: The act of assembling brawl, engaging in a brawl, resulting in the following consequences of serious injury.
-(4) Subjective Aspect: Direct intent, where the person knowingly and willfully engages in organizing or participating in the act of assembling brawl.
-Before discussing the Four-Element Theory (FET), it is necessary to briefly compare it with another key theory in Chinese criminal law, the Hierarchical Theory of Crime ConstitutionZhou (2017b); Zhang (2010), and the main distinction between these theories lies in whether a hierarchical structure is considered, with ongoing debates in practiceGao (2009); Chen (2010, 2017); Zhou (2017a).
-We chose FET as our foundational template for following reasons: 1) its dominance in Chinese judicial practice aligns with real-world criminal judgments; (2) its clear distinction between objective aspects and subjective intent offers direct reasoning checkpoints compared to the Three-Tier Theory; (3) its four-element annotation is flexible and can be adapted to the Three-Tier Theory by prioritizing objective analysis before subjective evaluationLi (2006); Zhang (2017).
-Recent approaches have leveraged the FET framework to model expert reasoning.
+LegalDiscourse shows that LLMs often fail to capture when laws apply and to whom Spangher et al.
+(2024), while LegalBench demonstrates that even state-of-the-art models underperform on diverse reasoning-intensive legal tasks Guha et al.
+(2023).
+To further enhance model performance, particularly in tasks requiring complex legal reasoning, some studies draw inspiration from established legal reasoning paradigms.
+For example, introducing the legal syllogism for legal judgment prediction Jiang and Yang (2023); using the IRAC paradigm to guide LLMs in reasoning about compositional rules Servantez et al.
+(2024).
+Several works have drawn on the FET in the context of Chinese criminal law.
 For example, breaking down legal rules into FET-aligned components using automated planning techniques Yuan et al.
+(2024); employing model-generated FETs as minor premises in legal judgment analysis Deng et al.
 (2023).
-Employing model-generated four-element structures as minor premises in legal judgment analysis Deng et al.
-(2023).
-While these methods have demonstrated improved performance on downstream tasks, they generally assume that the LLMs inherently understand the FET, without systematically validating this assumption.
-Notably, prior research on criminal charge prediction An et al.
-(2022) suggests that the models may misinterpret key legal concepts and may not be sensitive enough to the subtle differences in fact descriptions of confusing charges, highlighting the need to incorporate expert annotations to support LLM reasoning.
-3 Dataset Construction
-3.1 Hierarchical Legal Interpretation System
-Figure 1: Hierarchical Legal Interpretation System base on legal source validity.
-The legal sources follow a hierarchical order of validity.
-Thick arrows indicate the primary level where a particular interpretive method is applied, while dashed arrows represent its supplementary use at that level.
-Annotating the four elements of crime constitution is essentially a process of legal interpretation, which can be analyzed in two aspects:
-(1) What law is being interpreted.
-This involves identifying the sources of law, including statutory provisions corresponding to a specific charge, their associated judicial interpretations, case precedents, and academic discourses.
-In legal studies, these sources are categorized based on their legal validity into formal sources (which carry legal forces in judgments) and informal sources (which serve as references without legal forces)Pound (1925); Watson (1982); Pound (1932).
-Articles and judicial interpretations are considered formal sources, whereas case precedents and academic discourses are regarded as informal sources under the Chinese legal systemZhang and Zhou (2007).
-(2) How the law is interpreted.
-This pertains to legal interpretation methods, including literal interpretation, systematic interpretation, purposive interpretation, etc.
-These methods follow a hierarchical order in legal reasoningSutherland (1891); Kim and Division (2008); Eig Larry (2014).
-Legal interpretation should begin with literal interpretation (textual analysis).
-If the intended meaning cannot be clearly derived from the article alone, systematic interpretation and purpose interpretation should be applied first.
-If ambiguity remains, historical interpretation and comparative law interpretation may be used to further clarify the legal meaning.
-The specific definition is in AppendixB.
-Based on these principles, our annotation follows a pyramid structure of Hierarchical Legal Interpretation System base on legal source validity.
-As shown in Figure 1, the system is divided into two parts: Legal Source and Legal Interpretation Methods.
-The main structure of legal source follows a hierarchical order of validity: Article → Judicial Interpretations → Guiding Cases → Academic Discourses, where various legal interpretation methods are applied across different levels.
-Thick arrows indicate the primary level at which a particular method is used, while thin arrows denote the cross applications.
-3.2 Hierarchical Annotation Path of Legal Sources
-Our Annotators are experts who have all passed the National Uniform Legal Profession Qualification Examination and are familiar with the Four-Element Theory.
-The entire annotation process took a total of 7 months and involved 4 rounds of annotation according to the validity of the legal source from high to low level.
-The First Level: Article.
-Legal elements can be seen as an interpretation and refinement of the statutory provisions corresponding to a particular crime.
-Using literal interpretation as the primary method, the statute is broken down based on its semantic meaning and common usage, ensuring that the interpretation does not extend beyond the possible meaning of the text: (1) linguistic analysis follows the subject-predicate-object structure of the provision.
-(2) To maintain consistency, terms are systematically classified and mapped(e.g: subjective aspect is classified as either intentional or negligent.
-)(3) Only when it is impossible to make an explicit inclusion or exclusion judgment for an element based on the rules of language use (neutral option field), other interpretation methods should be used.
-This initial phase takes almost 2 months.
-For example, in the crime of robbery, the object "public or private property" represents the protected legal interest.
-The phrase "forcibly seizing public or private property through violence, coercion, or other means" describes the objective aspect.
-Since no subject is specified, it is assumed to involve a general subject, and the adverbs "violence" and "coercion" indicate an intentional act.
-Preliminarily interpret ‘violence’ in the objective aspect as ‘Use of physical force or power’, but the specific forms and subjects of violence need further clarification.
-The Second Level: Judicial Interpretation.
-In the 3rd and 4th months, the second level of the hierarchical annotation path focuses on refining legal elements through judicial interpretation.
-The primary method used for interpreting these materials is systematic interpretation.
-This approach examines the position of the corresponding articles within the legal system by analyzing their placement within the structure of laws, including parts, chapters, sections, articles, clauses, and subclauses, as well as their relationship to other statutes and judicial interpretations.
-Additionally, other interpretative methods, such as sociological interpretation and teleological interpretation, are referenced based on judicial interpretations, related statutory provisions, or bar exam questions.The goal of this level is to clarify the legislative intent by considering the contextual relevance of each provision within the broader legal framework.
-For example, in the first level, the objective aspect of "violence" in the crime of robbery requires further clarification, specifically regarding whether violence must be directed exclusively at persons or could also apply to property.
-Article 289 of Chinese Criminal LawCongress (2017) stipulates that in cases of "smashing, looting, and robbing" committed by a group, the ringleaders shall be convicted of robbery if they destroy or seize public or private property.
-This provision demonstrates that violence against property can also constitute robbery under Chinese law.
-The Third Level: Guiding Cases.
-In the 5th to 6th month, purposive interpretation and sociological interpretation are applied to the guiding cases and landmark judgments from the Supreme Court.
-By examining the social significance of real-world cases, these methods bridge the subtle gap between abstract legal theory and practical cases.
-This approach enables dynamic adaptation and integration of empirical insights and emerging controversies within the dataset.
-For example, in Criminal Trial Reference Case No.159Zou (2002), the perpetrator lured the victim into a room, locked the door, and seized 170,000 RMB intended for a transaction.
-The court determined that although the detention did not endanger personal safety, it was sufficient to suppress the victim’s resistance, thus constituting "violence" in in the objective aspects of robbery.
-Another example is the "Molestation and Theft Case" Ma (2021), where the perpetrator bound the victim, committed molestation, and stole the victim’s phone.
-Since the ongoing molestation reinforced coercion, it constitutes a new act of violence.
-Thus, the annotation includes "molestation" as an additional method.
-The Fourth Level: Academic Discourses.
-In the 7th month, the final stage involves academic expansion.
-Academic controversies are introduced by employing multiple interpretive methods such as comparative law interpretation, purposive interpretation, and sociological interpretation.
-These methods include inserting conflict markers at key points of controversy, highlighting the distinctions between mainstream consensus and minority theories, while providing brief annotations of their legal reasoning.
-This approach ensures the extensibility and academic depth of the dataset.
-For example, regarding the crime of robbery, for the main view in China, Soviet Union, North Korea, and Japan explicitly holds that the violence must be severe enough to endanger the victim’s life or healthZhang (2007).
-But some scholars argue that any violence that can forcibly impact the victim’s body is sufficient to constitute violence in robbery, no need to endanger the victim’s life or healthYang (2010).
-4 Data Distribution
-Metric LLM Expert
-Mean Median Mean Median
-Avg. Length 115.43 - 472.53 -
-SB 23.12 27 51.64 17
-OB 15.86 15 36.01 25
-SA 28.00 30 42.38 21
-OA 48.45 45 342.5 230
-Table 1: Comparison of Legal Element Lengths: LLM vs. Expert. SB = Subject, OB = Object, SA = Subjective Aspect, OA = Objective Aspect.
-As shown in Table 1, we compare the length of legal elements between expert-annotated descriptions in JUREX-4E and LLM-generated outputs across 105 charges that overlap with the Lecard-V2 dataset Li et al.
-(2024c), which is one of the most comprehensive legal datasets, covering 184 criminal charges.
-We find that:
-(1) The average total length of expert annotations (472.53) is more than four times longer than that of LLM-generated outputs (115.43), indicating that the former include more detailed information.
-(2) The median difference between the Subject (SB), Object (OB), and Subjective Aspect (SA) is relatively small, as these elements are typically fixed.
-For example, the SB is often a general entity, and the SA is often intent or negligence.
-(3) The median and mean values for SB and SA in the expert annotations differ, especially for SB (17 v.s.
-51.64).
-This discrepancy arises because certain specialized charges may require more detailed explanations.
-For example, in the crime of copyright infringement, the definition of “work” under the subject element has 9 occasions.
-Detailed data distribution for each element is provided in Appendix A.
-(4) The main difference between Expert and LLM is in the Objective Aspect (342.5 v.s.
-48.45 in Mean).
-This is because the OA includes a range of factual elements describing the criminal behavior, such as the conduct, object, result, time, and location, which are most emphasized in legal provisions and are central to various legal interpretive theories.
-5 Human Evaluation
-We selected 6 complicated crimes in Chinese judicial practiceOuyang et al.
-(1999) to evaluate whether the LLM can handle the Four-element Theory.
-Drawing from previous workDeng et al.
+While these methods have demonstrated improved performance on downstream tasks, they generally assume that the LLMs inherently understand the Four-Element Theory, without systematically validating this assumption.
+4 Can LLMs Grasp Legal Theory?
+To examine whether LLMs can understand and apply the Four-Element Theory (FET), we ask them to generate the four elements (FETs) for several representative charges and then analyze the outputs against expert annotations.
+We select GPT-4o as the target LLM, as it achieves state-of-the-art performance on open-source legal benchmarks Fei et al.
+(2023); Li et al.
+(2024c) and also outperforms others in our pilot study (Appendix D), indicating a strong capacity to understand and apply legal knowledge.
+Following prior work Deng et al.
 (2023); Cui et al.
 (2024); Zhou et al.
-(2023), we define LLM-generated knowledge as information produced by the LLM based on its pre-trained knowledge and contextual prompts.
-For detail, we provide the LLM with legal articles and the definition of each element in FET, prompting it to generate the four-elements base on these metrical.
-The LLM is expected to autonomously identify and generate the four elements based on its learned understanding of legal concepts.
-We invite legal experts to assess the four elements generated by the LLM from four dimensions: Precision, Completeness, Representativeness, and Standardization.
-  • Precision: Whether the key components of each element are accurately identified. This dimension mainly evaluates whether the four elements faithfully represent the legal provisions.
-  • Completeness: Whether all necessary information of each element is included.
-This assesses whether any essential content is missing, such as the omission of a description for specific subjects, like government officials.
-  • Representativeness: Whether the annotations highlight the most critical scenarios in judicial practice.
-For example, in crimes of intentional injury, this would involve describing the representative means of harm.
-  • Standardization: Whether the four elements are clearly defined, ensuring consistency in the expression of identical elements across different crimes (e.g., consistent description of general subjects), with concise and easily understandable explanations, free from legal ambiguities or misunderstandings.
-Each dimension was scored by two types of experts: one group with a pure legal background and another group with a combined background in law and Artificial Intelligence, all of whom have passed the bar examination.
+(2023), for each charge, we prompt GPT-4o with corresponding criminal articles (see prompt template in Appendix C).
+We invite legal experts who passed the National Judicial Examination to analyze the LLM-generated FETs and identify two main issues:
+(1) Inaccurate elements: LLMs may produce inaccurate FETs.
+For example, in Figure 1, for misappropriation of public funds, the LLM-generated Object is “the management order of public funds and the integrity of officials’ conduct”, missing the right to benefit from the use of public funds, which is necessary to identify this charge.
+(2) Insufficient interpretive ability: LLMs fail to recognize when statutory language requires deeper interpretation.
+As shown in Figure 1, the model simply extracts “misappropriating public funds for personal use” to describe the Objective Aspect.
+However, this phrase is far too general for practice.
+In judicial interpretations [^2] National People’s Congress Standing Committee.
+Interpretation on Article 384, Paragraph 1 of the Criminal Law of the People’s Republic of China, adopted at the 27th Meeting of the Standing Committee of the 9th National People’s Congress on April 28, 2002., the term “for personal use” should be interpreted with three situations: (1) using public funds for oneself, relatives, or other individuals; (2) lending public funds to other entities in one’s own name; or (3) using public funds in the name of one’s organization for another entity to gain personal benefits.
+5 Dataset Construction
+The lack of accuracy and interpretation in the generated FETs undermines the reliability of legal reasoning tasks.
+To address this, we introduce an expert-annotated FET dataset that captures both formal legal definitions and practical interpretive nuances, supporting more trustworthy and adaptable legal AI systems.
+To ensure both legal validity and interpretive clarity, we design a hierarchical annotation framework rooted in statutory sources and authoritative interpretive methods.
+5.1 Hierarchical Legal Interpretation System
+Figure 2: Hierarchical Legal Interpretation System based on legal source validity.
+The system consists of four annotation rounds, each using different interpretive methods based on different legal sources.
+Solid arrows indicate the primary method applied; dashed arrows represent supplementary use.
+Given a specific charge, we ask five legal experts to annotate the FETs based on relevant legal materials like articles and cases.
+This annotation process is essentially an act of legal interpretation.
+Legal interpretation refers to the application of various methods to analyze and understand legal texts, to determine their meaning and application in specific legal contexts Barak (2005).
+In our task, it involves applying different interpretation methods to the different materials in order to analyze and define the connotation and extension of each of the FETs of a charge.
+In designing our annotation framework, we address the following two questions:
+(1) What sources are interpreted.
+Legal interpretation draws upon various legal sources with different levels of validity.
+In legal studies, these sources are categorized based on their legal validity into formal sources (which carry legal force in judgments) and informal sources (which serve as references without legal force) Pound (1925); Watson (1982); Pound (1932).
+Articles and judicial interpretations are considered formal sources, whereas case precedents and academic discourses are regarded as informal sources under the Chinese legal system Zhang and Zhou (2007).
+Accordingly, we organize legal sources by their level of validity, with the following order of priority: Article → Judicial Interpretations → Guiding Cases → Academic Discourses.
+(2) How the law is interpreted.
+When interpreting the above sources, different interpretation methods are required.
+These methods follow a hierarchical order Sutherland (1891); Kim and Division (2008); Eig (2014): Legal interpretation should begin with literal interpretation (interpreting the text based on its plain meaning).
+If the intended meaning cannot be clearly derived from the article alone, systematic interpretation (considering the article’s role within the legal system) and purposive interpretation (considering the legislative intent) should be applied.
+If ambiguity remains, historical interpretation (based on the legislative history), sociological interpretation (based on the article’s social function and consequences), and other interpretation methods may be used to further clarify the legal meaning.
+The specific definition of legal interpretation methods is in Appendix B.
+We also consider the nature of each source.
+For example, Guiding Cases, as informal sources, do not define elements literally but instead supplement statutory interpretation through purposive, sociological, and other interpretive methods.
+The correspondence between interpretation methods and legal sources is illustrated by the orange arrows in Figure 2.
+5.2 Annotation Process
+As shown in the left part of Figure 2, our annotation process takes charges as input and outputs corresponding FETs, following a Hierarchical Legal Interpretation System to organize legal sources by validity and apply interpretation methods.
+The annotators are five experts, all of whom have passed the National Judicial Examination and are familiar with FET.
+The entire annotation process took 7 months and involved four rounds.
+Stage One: Literal interpretation using the core article.
+The interpretation of the FETs begins with the core article of each charge, which carries the highest legal validity, mainly through literal interpretation.
+At this stage, annotators analyze the article’s subject–predicate–object structure to identify candidate FETs, mapping the subject to the Subject (who commits the crime), the predicate (verb phrase) to the Objective Aspect (the conduct carried out), the object to the Object (the legal interest infringed), and adverbial phrases to the Subjective Aspect (the offender’s mental state).
+For example, Article 263 of the Chinese Criminal Law, concerning robbery, states that “forcibly seizing public or private property through violence, coercion, or other means” describes the Objective Aspect.
+Since the article lacks an explicit subject, a general subject is assumed by default.
+The adverbs “violence” and “coercion” indicate an intentional act.
+This stage spans two months.
+Stage Two: Systematic interpretation using related articles and judicial interpretations.
+While Stage One relies on the primary article of each charge, literal interpretation often leaves FETs underspecified.
+Stage Two, therefore, applies systematic interpretation, situating underspecified terms within the broader legal framework.
+By considering the provision’s function in the Criminal Law and its links to related articles and judicial interpretations (broadly understood here to include SPC guiding opinions and other interpretative documents), annotators clarify the scope and meaning of the element.
+For example, Article 263 of the Criminal Law does not specify whether “violence” must be directed only at persons or may also extend to property, nor does it make clear which borderline conduct should be excluded.
+Article 289 Congress (2017) provides that in mass “smashing, looting, and robbing,” the destruction or seizure of property by ringleaders may be punished as robbery, suggesting that violence may also cover acts against property.
+Conversely, the SPC’s 2016 Guiding Opinion on Robbery spc (2016) clarifies that if an offender uses only minor force to escape after theft, fraud, or snatching, and no injury above the statutory threshold results, such conduct is not deemed “violence” and does not requalify the offense as robbery.
+Stage Three: Purposive and sociological interpretation using guiding cases.
+Although the first two stages define the FETs based on articles and judicial interpretations, these sources remain abstract.
+In legal practice, courts also refer to Guiding Cases, designated by the Supreme People’s Court since 2011, to illustrate how legal articles are applied in concrete cases and interpreted in light of social purposes Chen et al.
+(2024).
+At this stage, annotators refine FETs by incorporating specific case scenarios from Guiding Cases.
+Since the number of Guiding Cases is limited, for rarer charges, we also consult model cases from the People’s Court Case Database and Gazette cases333retrieved from PKULaw: https://www.pkulaw.com/case?way=topGuid.
+Annotators mainly apply purposive and sociological interpretation to examine how legal elements are concretized in the reasoning process of practical cases, considering both legislative intent and social context.
+For example, in defining how “violence” in robbery operates in practice, annotators refer to a case involving molestation Ma (2021).
+The offender bound the victim to commit indecent acts and then took her phone.
+Under purposive and sociological interpretation, the ongoing molestation maintained the victim’s restrained state and thus constituted new violence.
+Such cases clarify that acts like molestation, though not physically injurious, can suppress resistance and therefore qualify as violence.
+This stage spans two months.
+Stage Four: Diverse interpretations using academic discourses.
+Although Stages One to Three refine both the abstract definitions and concrete scenarios of each element from various legal sources, certain elements still involve unresolved issues that rarely appear in practice and therefore lack clear judicial standards.
+At this stage, annotators consult academic discourses and apply diverse interpretations such as comparative, purposive, and sociological interpretation.
+For elements where disagreement exists, they record both mainstream and minority views, providing concise annotations that explain the underlying legal reasoning.
+For example, in defining “violence” in robbery, mainstream views in China, the former Soviet Union, North Korea, and Japan require that it endanger the victim’s life or health Zhang (2007b), while others argue that any force sufficient to subdue the victim should qualify Yang (2010).
+The annotations record both the dominant consensus and minority positions.
+This stage spans one month.
+The four stages represent the main interpretive approaches, but they are not mutually exclusive.
+In practice, annotators often combine methods when clarifying a particular element.
+As shown in Figure 2, the dashed orange arrows mark cross-stage interactions where multiple interpretive methods operate in a complementary way.
+5.3 Data Statistics
+Metric LLMMean LLMMedian ExpertMean ExpertMedian
+Avg. Length 115.43 - 472.53 -
+Subject 23.12 27 51.64 17
+Object 15.86 15 36.01 25
+Subjective Aspect 28.00 30 42.38 21
+Objective Aspect 48.45 45 342.5 230
+Table 1: Comparison of element lengths.
+Our dataset comprises 155 common criminal charges.
+These charges are selected based on their frequency in over 2.6 million publicly available criminal cases in China: specifically, we include all charges that appear more than 3,000 times, which together account for 91.71% of all cases (Appendix A), ensuring coverage of the most common real-world judicial scenarios.
+To compare the quality of expert-annotated FETs (expert-FETs) and LLM-generated FETs, we selected 105 charges in JUREX-4E that also appear in the widely used LeCaRDv2 dataset Li et al.
+(2024d).
+LLM-generated FETs were produced using the same setup as before, with a maximum output of 8192 tokens.
+Table 1 summarizes the differences in element length, with full length distributions available in Appendix A.
+Overall, expert-FETs are significantly longer, with an average total length of 472.53 words compared to 115.43 for LLM-generated ones.
+The most pronounced gap appears in the Objective Aspect (OA) (mean: 342.5 vs.
+48.45), where experts provide detailed factual descriptions, such as action, result, time, and location, often underdeveloped in LLM outputs.
+While the Subject (SB), Object (OB), and Subjective Aspect (SA) show smaller median differences, notable variation remains, especially in SB (mean: 51.64 vs.
+17), which in certain charges involves complex legal interpretations (e.g., “work” in copyright infringement) requiring more elaborate legal definitions.
+6 Human Evaluation
+To compare the quality of expert-annotated and LLM-generated FETs, we selected six complicated charges in Chinese judicial practice Ouyang et al.
+(1999).
+Based on prior theoretical framework Zhang (2007a), we assess the quality of FETs along four independent dimensions: Precision, Completeness, Representativeness, and Standardization.
+  • Precision: Whether each element accurately aligns with its statutory definition, reflecting key terms in the corresponding legal article.
+  • Completeness: Whether each element includes all practically necessary information, ensuring the definition is sufficient to guide legal reasoning.
+  • Representativeness: Whether the annotations reflect the most typical and practically significant scenarios in judicial practice.
+  • Standardization: Whether the expressions of elements are consistent across different charges, with clear, concise, and unambiguous language that facilitates understanding and minimizes interpretive variance.
+Evaluation was conducted by experts from two backgrounds: one group with a purely legal background and another with a combined background in law and AI, all of whom have passed the National Judicial Examination.
 The experts were selected to balance domain expertise and interdisciplinary perspectives.
 Scores were averaged across the two groups.
-Details about 1-5 scale criteria and annotator background are provided in Appendix C.
+Details about the 1-5 scale criteria and annotator background are provided in Appendix E.
 Dimension LLM Expert $\delta$
@@
 Standardization 4.33 4.56 + 0.23
-Table 2: Performance comparison of four elements across methods.
-$\delta$ represents the score difference between expert and LLM-generated four-elements, with experts outperforming LLMs in all dimensions.
-As shown in Table 2, expert annotations consistently outperform LLM-generated elements across all four dimensions, highlighting the limitations of LLMs in understanding legal elements.
+Table 2: Performance comparison of four elements across methods. $\delta$ represents the score difference between expert and LLM-generated FETs, with experts outperforming LLMs in all dimensions.
+As shown in Table 2, expert annotations consistently outperform LLM-generated elements across all four dimensions.
 The most pronounced deficiencies are observed in Completeness (+0.86) and Representativeness (+0.88).
-This suggests that while LLMs can generate formally standardized and relatively accurate four elements, their description are not specific enough and do not adequately reflect the representative features of a charge’s criminal composition.
-6 Evaluate Expert Knowledge on Charge Disambiguation
-In the preceding section, the human evaluation demonstrated that experts annotated higher-quality four-elements.
-To further quantitatively evaluate the annotations, a direct way is to judge whether different charges can be distinguished according to the four-element definition of crime constitution.
-Therefore, we introduce the Similar Charge Disambiguation (SCD) taskYuan et al.
+This aligns with our earlier analyses, where expert-generated elements include more factual details and representative descriptions.
+The gap in Precision (+0.57) suggests a tendency toward vague or legally irrelevant content, while the smaller difference in Standardization (+0.23) shows that LLMs can mimic structural patterns but lack deeper normative consistency.
+These results demonstrate the importance of expert supervision in providing reliable legal knowledge.
+7 Evaluation on Similar Charge Disambiguation
+To further validate annotation quality, we introduce the Similar Charge Disambiguation (SCD) task Yuan et al.
 (2024); Li et al.
 (2024a).
-6.1 Experiment Settings
-6.1.1 Dataset
-We chose the dataset released by Liu et al.
-(2021), which includes five charge sets with the largest number of cases.
-To evaluate performance on representative tasks, we selected three 2-label classification groups commonly examined in other datasets Yuan et al.
-(2024): Fraud & Extortion (F&E), Embezzlement & Misappropriation of Public Funds (E&MPF), and Abuse of Power & Dereliction of Duty (AP&DD).
-Each crime has over 1.9k cases, with a total of 13,962 cases.
-The details of the classification groups are shown in Appendix D.
-Following previous work Liu et al.
-(2021); Yuan et al.
-(2024), we use Average Accuracy (Acc) and macro-F1 (F1) as evaluation metrics.
-6.1.2 Baselines and Methods
-To evaluate SCD tasks, we consider two ways of incorporating legal knowledge.
-The first directly integrates legal statutes, represented by GPT-4o Achiam et al.
-(2023) as the baseline and GPT-4o+Article, which explicitly provides relevant legal articles to the model.
-The second adopts structured legal reasoning to enhance interpretability and accuracy.
-We consider Legal-CoT, a Chain-of-Thought Kojima et al.
-(2022) variant that conducts a stepwise analysis based on the FET, and MALR Yuan et al.
-(2024), a multi-agent framework that decomposes legal tasks into sub-tasks in four-element structures.
-Details of each baseline are provided in Appendix 7.
-We use an unified approach to introduce four-element descriptions.
-For each group of similar charges, the model receives charges’ four-elements from JUREX-4E or generated by LLM to aid classification.
-Specifically, GPT-4o+FETExpert relies on expert-annotated four-elements, while GPT-4o+FETLLM relies on LLM-generated four-elements.
-As shown in Appendix D, the instruction format is consistent across methods, with only the [Four Elements of candidate charges] varying based on the source.
-All experiments are conducted in a zero-shot setting, with the max_tokens set to 3,000 (or 10,000 for COT and MALR reasoning) and temperature set to 0 or 0.0001(In repeated experiments).
-6.2 Results
-Model F&E E&MPF AP&DD Average
+Given the case fact and a set of similar charges, SCD task requires the model to identify which charge is correct.
+We evaluate whether similar charges can be effectively distinguished based on their FETs, and whether expert-annotated FETs perform better than LLM-generated FETs.
+Model F&E E&MPF AP&$D$ Average
 Acc F1 Acc F1 Acc F1 Acc F1
@@
 MALR 94.62 95.82 86.99 86.98 87.86 88.68 89.82 90.49
-GPT-4o+FETLLM 95.73 96.56 91.87 92.01 89.61 89.69 92.40 92.75
+Farui-plus+FET4o 89.09 90.27 86.32 88.00 75.90 77.67 83.77 85.31
+Farui-plus+FETExpert 89.29 90.98 86.13 87.54 76.25 78.12 83.89 85.55
+Qwen2.5-72b+FET4o 93.15 95.06 90.99 93.56 87.71 88.56 90.62 92.39
+Qwen2.5-72b+FETExpert 93.29 95.18 91.18 93.66 87.81 89.45 90.76 92.76
+GPT-4o+FETfarui 94.86 96.12 91.84 92.64 89.35 89.85 92.02 92.87
+GPT-4o+FETqwen 95.53 96.53 91.82 92.96 89.48 90.09 92.28 93.19
+GPT-4o+FET4o+farui+qwen 94.97 96.24 91.84 92.73 89.69 90.12 92.17 93.03
+GPT-4o+FET4o 95.73 96.56 91.87 92.01 89.61 89.69 92.40 92.75
+GPT-4o+FET4o + ICL 95.74 96.36 91.84 92.01 90.48 90.63 92.69 93.00
 GPT-4o+FETExpert 96.06 96.69 92.57 93.05 90.53 90.62 93.05 93.45
-Table 3: Results of Charge Disambiguation. FET means introducing the Four-element theory with knowledge obtained from experts and LLM method. Highest results are in bold.
-As shown in Table 3, the GPT-4o+FETExpert performs best in discriminating similar charges, indicating that expert annotation is superior to other methods of directly or indirectly introducing FET with LLMs.
-Specifically, we can derive the following observation:
-Effectiveness of Domain-Specific Legal Knowledge:
-Among all approaches, those that explicitly incorporate domain-specific legal knowledge, such as GPT-4o+Article, Legal-CoT, and MALR, outperform GPT-4o alone.
-This highlights the importance of integrating legal knowledge.
-Importance of Concrete Four-element Knowledge:
-The accuracy of both Legal-CoT and MALR is still lower than GPT-4o+FET methods.
-This suggests that, compared to embedding the Four-Element Theory into LLMs’ reasoning process, providing concrete charge four-elements enables the model to better understand the different crimes’ composition.
-Superiority of Expert Annotations:
-Compared with the indirect introduction of FET reasoning, the method of directly introducing four-elements to the model (GPT-4o+FET) achieves better results.
-Notably, GPT-4o+FETExpert surpassing the GPT-4o+FETLLM by 0.65 in average accuracy and 0.70 in average F1-score, underscoring the superior quality and reliability of expert annotations in legal tasks, aligning with human evaluations in Table 2 and reaffirming the critical role of human expertise in legal decision-making.
-7 Can Expert Knowledge Benefit More Downstream Tasks?
-In this section, we design a simple framework to apply the expert-annotated four elements to Legal Case Retrieval (LCR), a task in which relevant cases are retrieved based on given facts.
-It is an important step in the practice of analyzing cases and making judgments, and it requires the precise application of the four-element theory to matches cases with similar criminal compositions.
-Figure 2: An expert-guided FET method to enhance legal case retrieval by incorporating expert four elements.
-7.1 Method
+Table 3: Performance on the Similar Charge Disambiguation (SCD) task.
+“Expert” refers to our expert-annotated FET, while “4o”, “qwen”, and “farui” refer to FET generated by different LLMs.
+Highest results are in bold.
+7.1 Experiment Settings
+7.1.1 Dataset and metrics
+We chose the SCD dataset released by Liu et al.
+(2021).
+Following previous work Yuan et al.
+(2024), we selected three 2-label classification groups: Fraud & Extortion (F&E), Embezzlement & Misappropriation of Public Funds (E&MPF), and Abuse of Power & Dereliction of Duty (AP&DD).
+Each charge has over 1.9k cases, with a total of 13,962 cases.
+The details of the groups are shown in Appendix F.
+Following previous work Liu et al.
+(2021); Yuan et al.
+(2024), we use Average Accuracy (Acc) and macro-F1 (F1) as evaluation metrics.
+7.1.2 Baselines and Methods
+We compared the following baselines: GPT-4o Achiam et al.
+(2023) and GPT-4o+Article, which explicitly supplies relevant legal articles; Legal-COT Kojima et al.
+(2022), a Chain-of-Thought variant that applies the Four-Element Theory step by step, and MALR Yuan et al.
+(2024), a multi-agent framework that decomposes legal tasks into FET-aligned subtasks.
+Details are in Appendix 9.
+Methods: Following Section 4, our main model is GPT-4o.
+We also compared Farui-plus (the latest version of Tongyifarui, representative legal LLM) and Qwen2.5-72B Bai et al.
+(2023) (representative open-source LLM).
+To incorporate FET knowledge, each group of similar charges is augmented with four-element descriptions, either generated by LLMs or sourced from JUREX-4E.
+For example, GPT-4o+FETLLM uses LLM-generated FETs, while GPT-4o+FETExpert uses expert-annotated ones.
+The input format is fixed across methods, differing only in the [four-elements of candidate charges] (Appendix F).
+All experiments are zero-shot, with max_tokens set to 3,000 (10,000 for Legal-COT and MALR) and a temperature of 0 or 0.0001 in repeated runs.
+To further explore generation setups, we also evaluate an ICL variant, GPT-4o+FET4o + ICL, where two representative FET exemplars (Theft and Snatching) are provided in the prompt to guide generation.
+7.2 Results
+The SCD results are shown in Table 3, where we can observe that:
+Effectiveness of Structured FET Knowledge: Providing specific structured charge FETs yields the highest accuracy among all legal knowledge integration methods.
+Compared to implicit approaches, such as prompts (GPT-4o+Article, Acc 92.09) or reasoning chains (Legal-COT, Acc 89.95), structured FET knowledge offers more effective support for legal decision-making (e.g., GPT-4o+FETExpert, Acc 93.05)
+Superiority of Expert-Annotated FET: Expert-annotated FET consistently outperforms LLM-generated FET across three representative LLMs, including FETfarui, FETqwen, FET4o, and their combination (FET4o+farui+qwen).
+For example, GPT-4o+FETExpert surpasses GPT-4o+FET4o by 0.65 in average accuracy and 0.70 in F1-score.
+Consistent Gains Across Models: Expert-annotated FETs yield consistent performance gains across different SCD models.
+When applied to Farui-plus, Qwen2.5-72b, and GPT-4o, it improves F1-score by +0.24, +0.37, and +0.70, respectively over their LLM-generated FET baselines.
+The ICL-based variant yields consistent improvements over direct prompting (Table 3), demonstrating the benefit of exemplar guidance, though it still lags behind expert-FETs.
+We also conducted McNemar’s test on paired samples between each LLM-generated FET and the expert-FET, which shows statistically significant improvements ($p<0.05$) across all tasks (see Table 11 in Appendix F).
+8 Application in Legal Case Retrieval
+In this section, we design a simple expert-guided FET method to apply JUREX-4E to the Legal Case Retrieval (LCR) task, which retrieves relevant cases based on case facts.
+This task is well-suited for FET because it requires a comprehensive comparison of the four elements across different charges in cases.
 Model NDCG@10 NDCG@20 NDCG@30 R@1 R@5 R@10 R@20 R@30 MRR
-BERT 0.1511 0.1794 0.1978 0.0199 0.0753 0.1299 0.2157 0.2579 0.1136
-Legal-BERT 0.1300 0.1487 0.1649 0.0186 0.0542 0.1309 0.1822 0.2172 0.0573
-Lawformer 0.2684 0.3049 0.3560 0.0432 0.1479 0.2330 0.3349 0.4683 0.1096
-ChatLaw 0.2049 0.2328 0.2745 0.0353 0.1306 0.1913 0.2684 0.3751 0.1285
-SAILER 0.3142 0.4133 0.4745 0.0539 0.1780 0.3442 0.5688 0.7092 0.1427
-GEAR * * * 0.0630 0.1706 0.3142 0.4625 * 0.2162
-BGE 0.4737 0.5539 0.5937 0.0793 0.2945 0.4298 0.6500 0.7394 0.1926
-FETLLM 0.5139 0.5862 0.6291 0.0980 0.2967 0.4769 0.6802 0.7828 0.2140
-        - base 0.3583 0.4293 0.4798 0.0506 0.2240 0.3644 0.5383 0.6652 0.1453
-FETExpert_guided 0.5211 0.5920 0.6379 0.1024 0.3049 0.4883 0.6885 0.7967 0.2155
-        - base 0.3766 0.4584 0.5111 0.0715 0.1894 0.3709 0.5891 0.7203 0.1624
-Table 4: SCR results. Bold fonts indicate leading results in each setting. * denotes that the indicator is not applicable to the current model.
-We implement a standard dense retrieval approach BGE using BGE-m3 Chen et al.
-(2023), an advanced embedding model for dense retrieval.
-Given a query $q$ and a candidate case $c$, their vector representations $\mathbf{v}_{q}$ and $\mathbf{v}_{c}$ are obtained through shared encoder $E$: $\mathbf{v}_{q}=E(q),\quad\mathbf{v}_{c}=E(c)$.
-We used the BGE-m3 model without fine-tuning as the shared encoder.
-Next, the relevance score is computed via cosine similarity:
-$\text{sim}_{\text{base}}(q,c)=\frac{\mathbf{v}_{q}\cdot\mathbf{v}_{c}}{\|\mathbf{v}_{q}\|\|\mathbf{v}_{c}\|}$ (1)
-To retrieve the top-k most similar cases, we rank the candidates based on their cosine similarity to the query.
-Denote the set of candidate cases as $C=\{c_{1},c_{2},\dots,c_{n}\}$, where $n$ is the total number of candidate cases.
-We compute the similarity for each $c_{i}\in C$, and select the top-k candidates with the highest similarity scores.
-As shown in Figure 2, to leverages expert-annotated four elements of charges, we introduce an BGE+FETExpert_guided method for the retrieval process, consisting of three steps: (1) Predicting charges, a LLM $\mathcal{M}_{p}$ predicts potential charges $Z=\{z_{1},...,z_{k}\}$ from case facts.
-(2) Matching elements, retrieving corresponding charge’s four-elements $\{f_{z}\}_{z\in Z}$ in JUREX-4E.
-(3) Analyzing case facts.
-Guided by $\{f_{z}\}$, another LLM $\mathcal{M}_{g}$ generates case-specific four elements $a_{c}$ for candidate $c$.
-The final similarity score combines factual and theoretical alignment:
-$\text{sim}_{\text{final}}(q,c)=\alpha\cdot\text{sim}_{\text{base}}(q,c)+(1-\alpha)\cdot\text{sim}_{\text{f}}(a_{q},a_{c})$ (2)
-where $\alpha=0.7$ and $\text{sim}_{\text{f}}$ measures the similarity between the generated four-element descriptions.
-To facilitate comparison, we also design a BGE+FETLLM method that directly prompt the LLM $\mathcal{M}_{g}$ with the concept of Four-Element Theory to generate case-specific four elements $a_{c}$.
-7.2 Dataset
+BGE (case_fact only) 0.4737 0.5539 0.5937 0.0793 0.2945 0.4298 0.6500 0.7394 0.1926
+BGE+FET (Qwen2.5) 0.5125 0.5858 0.6350 0.1104 0.2870 0.4653 0.6679 0.7836 0.2168
+   FET only 0.3367 0.3971 0.4487 0.0622 0.2006 0.3279 0.4806 0.6037 0.1524
+BGE+FET (Expert, Qwen2.5) 0.5295 0.5979 0.6416 0.1124 0.3122 0.4838 0.6791 0.7824 0.2206
+   FET only 0.3354 0.4035 0.4541 0.0849 0.1923 0.3076 0.4839 0.6097 0.1606
+BGE+FET (GPT-4o) 0.5139 0.5862 0.6291 0.0980 0.2967 0.4769 0.6802 0.7828 0.2140
+   FET only 0.3583 0.4293 0.4798 0.0506 0.2240 0.3644 0.5383 0.6652 0.1453
+BGE+FET (Expert, GPT-4o) 0.5211 0.5920 0.6379 0.1024 0.3049 0.4883 0.6885 0.7967 0.2155
+   FET only 0.3766 0.4584 0.5111 0.0715 0.1894 0.3709 0.5891 0.7203 0.1624
+Table 4: Performance on the Legal Case Retrieval (LCR) task. The highest results are in bold. “FET only” indicates using the four-element descriptions without case facts.
+8.1 Dataset and Metrics
 LeCaRDv2 Li et al.
-(2024c) is the latest version of LeCaRDMa et al.
-(2021), which is widely used in LCR task Li et al.
+(2024d) is the latest version of LeCaRD Ma et al.
+(2021), which is widely used in legal tasks Li et al.
 (2024b); Zhou et al.
@@
 It comprises 800 queries and 55,192 candidates extracted from 4.3 million criminal case documents.
-There are two common evaluation settings for this dataset: one uses a subset Qin et al.
-(2024) with a candidate pool size of 1,390, while the other uses the full set Li et al.
-(2024c) with a candidate pool size of 55,000.
-We conducted experiments under both settings.
-Following previous workFeng et al.
-(2024); Qin et al.
-(2024), we adopt commonly used evaluation metrics.
-For the subset, we use NDCG@10, 20, 30, Recall@1, 5, 10, 20, and MRR.
-For the full dataset, we use Recall@100, Recall@200, Recall@500, and Recall@1000.
-7.3 Baselines
-Consistent with earlier workLi et al.
-(2024c); Qin et al.
-(2024), we compare some dense retrieval methods, including: BERTDevlin (2018), LawformerXiao et al.
-(2021), ChatLaw-Text2Vec [^1] https://modelscope.cn/models/fengshan/ChatLaw-Text2VecCui et al.
-(2023), SAILERLi et al.
-(2023), GEARQin et al.
-(2024).
-Details of each baseline is shown in Appendix E.
-These baselines are implemented using the FlagEmbedding Toolkit [^2] https://github.com/FlagOpen/FlagEmbedding with a RTX 3090.
+Following previous work Qin et al.
+(2024), we chose 1390 candidates and used NDCG@10, 20, 30, Recall@1, 5, 10, 20, and MRR as metrics.
+We also tested different candidate pool settings (see Appendix G).
+The results are consistent.
+8.2 Baselines and Methods
+We adopt a dense retrieval framework based on BGE-m3 Chen et al.
+(2023), a strong embedding model for legal and general-domain texts.
+Given a query $q$ and a candidate case $c$, we compute their vector representations using a shared BGE-m3 encoder.
+Retrieval is performed by computing cosine similarities between the query and all candidates and selecting the top-k candidates.
+To enhance retrieval accuracy, we compare the following three methods:
+(1) BGE(case_fact only): Standard dense retrieval using only BGE-m3 embeddings of the raw case facts.
+(2) BGE+FET ($\mathcal{M}_{g}$): We prompt different LLMs $\mathcal{M}_{g}$ to generate a structured four-element description of each case (case-FET) based solely on its facts, without using external knowledge.
+These case-FETs are then embedded with BGE-m3 and used to compute similarity.
+Because the FET abstracts away case-specific details, we combine the original fact-based similarity and the FET-based similarity in a ratio of 7:3.
+(3) BGE+FET (Expert, $\mathcal{M}_{g}$): An expert-guided FET method that incorporates JUREX-4E to guide case-FET generation. It consists of four steps:
+  1. 1.
+    Charge Prediction. A charge prediction model $\mathcal{M}_{p}$ (Qwen-plus, details see Appendix D) predicts the set of likely charges $Z=\{z_{1},...,z_{k}\}$ for the query case.
+  2. 2.
+    Expert-FET Matching. Retrieving corresponding charge’s four-elements $\{f_{z}\}_{z\in Z}$ for each predicted charge in JUREX-4E. These provide theoretical guidance for subsequent reasoning.
+  3. 3.
+    Case-FET Generation. Guided by $\{f_{z}\}$, the LLM $\mathcal{M}_{g}$ generates case-specific four-elements $fet_{c}$ for candidate $c$.
+  4. 4.
+    Dense retrieval. We embed the generated FETs using BGE-m3 and compute similarity scores as in Method (2), combining both factual and FET-based similarities.
+For the $\mathcal{M}_{g}$, we chose Qwen2.5-72b and GPT-4o.
+The retrieval framework is implemented with the FlagEmbedding Toolkit [^4] https://github.com/FlagOpen/FlagEmbedding with an RTX 3090.
+Following prior work Li et al.
+(2024d); Qin et al.
+(2024), we also compare some dense retrieval methods to examine the representativeness of BGE-m3.
+Results of baselines and prompt templates are available in Appendix G.
 8.3 Results
-The LCR results are shown in Table 4, where we can observe that:
-FET Works Well in LCR.
-The baseline model BGE achieves strong performance across most metrics compared to previous methods.
-Introducing the Four-Element Theory (FET) further improves its results, with relative MRR improvements of 11.11% for FETLLM and 11.89% for FETExpert_guided, indicating that introducing legal theory is effective.
-Expert Knowledge is Necessary.
-By leveraging external knowledge, FETExpert_guided achieves significant improvements across all of the metrics.
-Specifically, using expert-guided case four-elements (FETExpert_guided-base) outperforms LLM-generated case four-elements (FETLLM-base) by an average of 11.77% in MRR, demonstrating the critical role of expert knowledge in enhancing retrieval precision.
-A case study in Appendix G shows that the expert four-element for charges provide practical judgment points and key narratives (e.g., the special subject of the Crime of Embezzlement) that help the LLM focus on essential facts to analyze the case.
-We also evaluated the FET method on the full set, as shown in Table 9 , and the results remain consistent, with the expert-guided method still performing best.
+The LCR results are shown in Table 4, where we can observe that: (1) FET Enhances Retrieval.
+Integrating the FET improves retrieval performance across all metrics.
+For instance, BGE+FET(GPT-4o) improves MRR by 11.11%, and BGE+FET (Expert, GPT-4o) achieves an even larger gain of 11.89%, indicating that structured legal theory benefits retrieval quality.
+(2) Expert Knowledge is Important.
+Expert-guided case-FET consistently outperforms LLM-generated variants across both Qwen2.5-72b and GPT-4o backbones.
+For example, BGE+FET (Expert, GPT-4o) achieves higher Recall@30 (0.7967 vs.
+0.7828) and MRR (0.2155 vs.
+0.2140).
+The gap is even larger in the FET only setting (e.g., MRR 0.1624 vs.
+0.1453 for GPT-4o), demonstrating that expert knowledge captures critical legal reasoning that LLMs may overlook.
+We provide a case study in Appendix I.
+It illustrates that the expert-annotated FETs of charges provide practical judgment points and key narratives (e.g., the special subject of Embezzlement) that help the LLM focus on essential facts to analyze the case-FET.
+9 Discussion
+Model F&E Acc F&E F1 E&MPF Acc E&MPF F1 AP&DD Acc AP&DD F1 Avg. Acc Avg. F1
+GPT-4o+FET4o 95.73 96.56 91.87 92.01 89.61 89.69 92.40 92.75
+GPT-4o+FETCollaboration 95.99 96.56 91.51 91.61 91.02 90.95 92.84 93.04
+GPT-4o+FETExpert 96.06 96.69 92.57 93.05 90.53 90.62 93.05 93.45
+Table 5: Results of the collaboration study on Similar Charge Disambiguation (SCD).
+To examine whether LLMs can effectively support expert annotation, we designed a preliminary collaboration pipeline that is roughly aligned with our annotation stages.
+For each charge, the LLM first retrieved relevant statutory articles (the retrieval of legal sources in stages 1–2).
+It then extracted the top 10 factual keywords from cases cited by experts (the use of factual cues as in stage 3).
+Experts validated these keywords for each element, after which the LLM generated refined FET annotations by reasoning over the combined legal texts and factual cues.
+This pipeline yielded measurable improvements over unguided prompting (Table 5), illustrating the feasibility of expert-in-the-loop workflows.
+Nevertheless, the study also highlights limitations.
+High-frequency keywords often lacked discriminative power for similar charges—for example, descriptors such as “beating” and “stabbing” occurred in both intentional injury and intentional homicide, reducing their utility.
+The refined FETs still fell short of the depth and precision of expert annotations.
+Future improvements could include decomposing FET generation into subtasks aligned with annotation stages and enriching each stage with more authoritative sources, thereby strengthening both coverage and normative awareness.
 10 Conclusion
-In this paper, we propose an expert-annotated knowledge base, evaluate its quality in the Similar Charge Distinction task, and apply it to the Legal Case Retrieval task.
-Our results demonstrate that expert annotations significantly enhance LLMs’ understanding of the Four-Element Theory.
-The four-element annotations, enriched with professional legal interpretations, provide strong support for LLMs’ reasoning capabilities.
-This approach can be extended to other legal AI tasks, such as legal document analysis and contract interpretation.
+This paper presents JUREX-4E, an expert-annotated FET knowledge base built through a structured legal interpretation process and validated on downstream tasks.
+Grounded in widely accepted interpretive methods, our framework is adaptable across different branches of law and legal traditions, making it applicable beyond Chinese criminal law.
+Moreover, the structured approach to integrating expert domain knowledge may inspire applications in other areas where expert judgment is critical.
 11 Ethical Considerations
@@
 12 Limitations
-As a limitation, this knowledge base focuses on the Four-Element Theory within the context of 155 crimes under Chinese Criminal Law.
-However, the four-level hierarchical pyramid annotation structure based on the legal interpretation system proposed in this work provides valuable insights for future expansion to other legal domains, as it represents a theoretical framework in the field of jurisprudence.
-The interpretative methods within the legal interpretation system, including textual, systematic, sociological, and doctrinal interpretations, are universally recognized in international law field and can be applied to different laws, countries, and legal systems.
+Our current knowledge base is limited to 155 charges under Chinese Criminal Law due to the high cost of expert annotation.
+Future work will explore extending it to other legal domains and jurisdictions.
+Another limitation lies in our current integration of factual and legal information.
+In the LCR task, although case facts are used to generate FETs, the FET only variant excludes the original case facts during retrieval, resulting in performance loss (e.g., MRR 0.1624 vs.
+0.2155).
+This suggests that our current method remains coarse-grained, and more fine-grained fusion strategies, such as multi-agent coordination or retrieval-time integration, deserve future exploration.
 Acknowledgments
-Appendix A Detailed Data Distribution for each Element
-Figure 3: The average length distribution of total four elements annotated by experts.
+This work is supported in part by Beijing Science and Technology Program (Z231100007423011)
+Appendix A Charge Selection and Detailed Data Distribution
+Charge Selection: To systematically determine charge frequency, we analyzed the CAIL2018 dataset Xiao et al.
+(2018), which contains 2,676,075 criminal cases annotated with 183 criminal law articles and 202 criminal charges and a total of 3,010,000 criminal charges.
+Apart from few charges that have been merged or changed name, our dataset largely covers all criminal charges from CAIL2018 that have a frequency of over 3,000 (>0.099%) occurrences.
+Length distribution for each element: Table 6.
+Figure 3: The average length distribution of the total four elements annotated by experts.
 Figure 4: The length distribution of each element annotated by experts.
-Figure 5: The average length distribution of total four elements generated by LLM.
+Figure 5: The average length distribution of the total four elements generated by LLM.
 Figure 6: The length distribution of each element generated by LLM.
@@
 2. Systematic Interpretation
-An approach interpreting legal provisions through their position within the codified legal hierarchy and logical connections with related norms, maintaining the integrity of the legal system (aligned with Dworkin’s "law as integrity" theory).
+An approach interpreting legal articles through their position within the codified legal hierarchy and logical connections with related norms, maintaining the integrity of the legal system (aligned with Dworkin’s "law as integrity" theory).
 3. Purposive Interpretation
@@
 4. Historical Interpretation
-Interpretation based on legislative history materials including drafts, debates and official commentaries, while distinguishing original meaning from framers’ subjective intentions (as per Brest’s original understanding theory).
+Interpretation based on legislative history materials, including drafts, debates, and official commentaries, while distinguishing original meaning from framers’ subjective intentions (as per Brest’s original understanding theory).
 5. Comparative Interpretation
@@
 6. Sociological Interpretation
-Interpretation evaluating social efficacy through empirical analysis of implementation effects, guided by Pound’s sociological jurisprudence principle that "law must be measured by its achieved results".
-Appendix C Human Evaluation Guidance
+Interpretation evaluating social efficacy through empirical analysis of implementation effects, guided by Pound’s sociological jurisprudence principle that “law must be measured by its achieved results”.
+Appendix C Prompt for LLM-generated FET
+See Table 6.
+You are an expert in criminal law. Based on the given charge, please analyze it according to China’s criminal law and output the four elements of the charge in order, including:
+- Object: The concretization of a certain abstract social interest.
+For example, the object of charges that infringe on personal rights is the right to life, while the object of property-related charges could be items such as mobile phones or wallets.
+- Objective Aspect: The objective facts of the criminal act, including the key actions that trigger the charge (e.g., theft, robbery) and the consequences caused by the act (e.g., serious injury, death, property loss).
+- Subject: Typically, the general subject of the charge, but in some cases, a specific subject is required (e.g., government officials in certain offenses).
+- Subjective Aspect: The mental state of the perpetrator, such as intent or negligence.
+Relevant Legal Articles: []
+Please synthesize the above information to generate a refined set of four elements that represent the characteristics of the charge.
+Output format: { "Crime": "", "Four-elements of the Crime": { "Crime Object": "", "Objective Aspect": "", "Subject": "", "Subjective Aspect": "" } }
+Crime: []
+Table 6: Prompt template for generating the four elements of a Crime using LLMs
+Appendix D Details about Pilot Study
+We selected candidate models from LawBench Fei et al.
+(2023) and LexEval Li et al.
+(2024c), which contain the broadest and most up-to-date evaluation of legal LLMs.
+From these, we chose top-performing models such as GPT-4, Qwen-14B-chat, and representative legal-specific LLMs.
+For best performance, we used GPT-4o (the latest version of GPT-4 at that time) and Qwen-plus (a stronger commercial variant of Qwen2.5-72B. (Aliyun model-studio official site))
+During implementation, we found that most legal LLMs were unavailable.
+The only stably accessible one was Farui (A leading legal LLM built on Qwen, (Aliyun model-studio official site, Tongyi Farui)), specifically the version “tongyifarui-890” from its official API.
+To compare GPT-4o, Qwen-plus, and tongyifarui-890, we sampled 300 cases from our legal retrieval dataset and asked models to perform charge prediction, which is the pre-task for generating case-FETs.
+(For each case in legal retrieval, the model was required to predict charges, so we can match charges’ expert-FETs, and use them to generate case-FET.)
+This task involved all criminal charges, including multi-defendant and multi-charge scenarios, and requires models to predict charges from open text without a predefined list, making it a challenging legal task.
+The result showed that GPT-4o (59.78%) > Qwen-plus (58.70%) >> tongyifarui-890 (21%). Given Farui’s poor performance, we did not include it in subsequent experiments.
+We further evaluated GPT-4o and Qwen-plus based on their ability to generate case-FETs.
+The results showed that GPT-4o outperformed Qwen-plus (MRR 0.2140 vs.
+0.2052).
+Considering both results, we adopted GPT-4o as our primary model in the paper.
+Subsequently, in efforts to improve charge prediction for matching charges’ expert-FETs, we found that Qwen-plus performed better than GPT-4o when a charge list was provided (58.70%->80.43% vs.
+59.78%->71.74%).
+Therefore, in this specific setting for charge prediction before retrieval, we used Qwen-plus.
+For fair and reproducible presentation of results on specific downstream tasks (SCD and LCR), as mentioned in the main text, we present the results of open source Qwen2.5-72b.
+Appendix E Human Evaluation Guidance
 The annotators included three postgraduate students specializing in criminal law and one master’s student in legal science and technology.
@@
 Before scoring, they were asked to read the descriptions and scoring guidelines (as shown in Table 7) for each evaluation dimension.
-In order to ensure the fairness of the evaluation, they do not know the source of each four elements, and even do not know that these four elements include those generated by LLMs.
+In order to ensure the fairness of the evaluation, they do not know the source of the four elements, and they do not know that the four elements include those generated by LLMs.
 When assigning scores, they were also required to provide brief justifications.
@@
 Dimension Precision Completeness Representativeness Standardization
-Definition Whether there are errors in key elements Whether the four elements are complete Whether key elements and scenarios are emphasized Whether language and format are clear and standardized
+Definition Whether there are errors in key elements Whether the four-elements are complete Whether key elements and scenarios are emphasized Whether language and format are clear and standardized
 Score 1 Contains numerous obvious errors, severely impeding the judgment of culpability, exculpation, and conviction, leading to significant deviations.
@@
 Table 7: The four dimensions of the human evaluation and the specific score description.
-Appendix D Details for Similar Charge Disambiguation
+Appendix F Details for Similar Charge Disambiguation
 Charge Sets Charges Cases
@@
 E&MPF Embezzlement & Misappropriation of Public Funds 2391 / 1998
-AP&DD Abuse of Power & Dereliction of Duty 1950 / 1938
+AP&$D$ Abuse of Power & Dereliction of Duty 1950 / 1938
 Table 8: Distribution of charges in the GCI dataset.
-Cases denotes the number of cases in each category.
+Cases denote the number of cases in each category.
 Following Liu et al.
@@
 Please determine which of the following candidate charges the given facts align with.
-The candidate charges and their corresponding four elements are as follows:
-[Four Elements of Candidate Charges].
+The candidate charges and their corresponding four-elements are as follows:
+[four-elements of Candidate Charges].
 The four elements represent the core factors for determining the constitution of a criminal charge.
@@
 Please compare the case facts to determine which charge’s four elements they align with, thereby identifying the charge.
-Table 7: Prompt template for adding the Four-Element Theory and specific four elements of crime in charge disambiguation.
+Table 9: Prompt template for adding the Four-Element Theory and specific four-elements of crime in charge disambiguation.
 For LLM baselines, we evaluate both general-purpose and task-specific methods.
 GPT-4o is an optimized version of GPT-4 Achiam et al. (2023) that has well performance in specific tasks through domain adaptation.
-To explore the effectiveness of notes-guided four elements in LLMs, we further consider other methods that introduced the Four-element theory into LLMs.
+To explore the effectiveness of expert-FETs, we further consider other methods that introduced the Four-element Theory into LLMs.
 GPT-4oLaw, which introduces articles related to corresponding charges into the instruction to provide legal context.
 Legal-COT is a variant of COT Kojima et al. (2022) that guides the LLM to perform step-by-step legal reasoning by incorporating explanations of the Four-element theory into the instruction.
-MALR is a up to date multi-agent framework designed to enhance complex legal reasoning Yuan et al.
+MALR is an up-to-date multi-agent framework designed to enhance complex legal reasoning Yuan et al.
 (2024), enabling LLMs to autonomously decompose legal tasks and extract insights from legal rules.
@@
 Experiments on the paper’s reported examples show that our implementation produces task decompositions and outputs largely consistent with the original results.
-As shown in Table D, different methods differ in their prompts for generating and explaining the Four-Element Theory, but generally follow a similar process.
+As shown in Table 10, different methods differ in their prompts for generating and explaining the Four-Element Theory, but generally follow a similar process.
 For the SCD output, except for COT and MALR, which require reasoning processes and prediction results, all other methods only require the output of prediction results.
 Method GPT-4o GPT-4o+Article Legal-COT GPT-4o+FETLLM GPT-4o+FETExperts
-Pre-task None None None LLM-generated four elements Expert-annotated four elements
+Pre-task None None None LLM-generated FETs Expert-annotated FETs
 Prompt You are a lawyer specializing in criminal law. Based on Chinese criminal law, please determine which of the following candidate charges the given facts align with.
-Candidate charges are as follows: #Candidate Charges The candidate charges and relevant legal articles are as follows: #Candidate Charges + #Articles Please analyze using the Four Elements Theory step by step: #details about each step.
-The candidate charges are as follows: #Candidate Charges The candidate charges and their corresponding four elements are as follows: #Four Elements of candidate charges.
+Candidate charges are as follows: #Candidate Charges The candidate charges and relevant legal articles are as follows: #Candidate Charges + #Articles Please analyze using the four-elements Theory step by step: #details about each step.
+The candidate charges are as follows: #Candidate Charges The candidate charges and their corresponding four-elements are as follows: #four-elements of candidate charges.
 The four elements represent the four core factors of a charge.
 Compare the case facts to determine which charge’s four elements they align with, thereby identifying the charge.
-Output format: #Format. Note: Only output the charge, no additional information.
-Case facts: #Case Facts.
+Output format: #Format. Note: Only output the charge, no additional information. Case facts: #Case Facts.
 Table 10: Prompts of different methods in Similar Charge Disambiguation. # represents a format input.
-Appendix E Baselines in Legal Case Retrieval
+FET-LLM vs FET-Expert F&E E&MPF AP&$D$
+FET-qwen vs FET-Expert 0.00215 0.00070 0.00509
+FET-farui vs FET-Expert 0.00000 0.00126 0.00516
+FET-4o+farui+qwen vs FET-Expert 0.00000 0.00996 0.03415
+FET-4o vs FET-Expert 0.02246 0.00156 0.02251
+Table 11: McNemar’s test results (p-values) comparing LLM-generated FET and expert-FET. Statistically significant improvements ($p<0.05$) are observed across all tasks.
+Appendix G Baselines in Legal Case Retrieval
 BERT Devlin (2018) is a language model widely used in retrieval tasks.
@@
 (2023) is a structure-aware legal case retrieval model utilizing the structural information in legal case documents.
-GEARQin et al.
-(2024) is a generative retrieval framework that explicitly integrates judgment prediction with legal document retrieval in a sequence-to-sequence manner.
-Since the output of GEAR cannot directly evaluate NDCG, the official results under the same setting are directly referenced in this paper.
-LLM and Expert represent the results of retrieval using only the four elements.
-Appendix F SCR results on the full LeCaRDv2 Dataset
+Baseline results are provided in Table 12.
+Model NDCG@10 NDCG@20 NDCG@30 R@1 R@5 R@10 R@20 R@30 MRR
+BERT 0.1511 0.1794 0.1978 0.0199 0.0753 0.1299 0.2157 0.2579 0.1136
+Legal-BERT 0.1300 0.1487 0.1649 0.0186 0.0542 0.1309 0.1822 0.2172 0.0573
+Lawformer 0.2684 0.3049 0.3560 0.0432 0.1479 0.2330 0.3349 0.4683 0.1096
+ChatLaw-Text2Vec 0.2049 0.2328 0.2745 0.0353 0.1306 0.1913 0.2684 0.3751 0.1285
+SAILER 0.3142 0.4133 0.4745 0.0539 0.1780 0.3442 0.5688 0.7092 0.1427
+BGE (case_fact only) 0.4737 0.5539 0.5937 0.0793 0.2945 0.4298 0.6500 0.7394 0.1926
+BGE+FET (Qwen2.5) 0.5125 0.5858 0.6350 0.1104 0.2870 0.4653 0.6679 0.7836 0.2168
+   FET only 0.3367 0.3971 0.4487 0.0622 0.2006 0.3279 0.4806 0.6037 0.1524
+BGE+FET (Expert, Qwen2.5) 0.5295 0.5979 0.6416 0.1124 0.3122 0.4838 0.6791 0.7824 0.2206
+   FET only 0.3354 0.4035 0.4541 0.0849 0.1923 0.3076 0.4839 0.6097 0.1606
+BGE+FET (GPT-4o) 0.5139 0.5862 0.6291 0.0980 0.2967 0.4769 0.6802 0.7828 0.2140
+   FET only 0.3583 0.4293 0.4798 0.0506 0.2240 0.3644 0.5383 0.6652 0.1453
+BGE+FET (Expert, GPT-4o) 0.5211 0.5920 0.6379 0.1024 0.3049 0.4883 0.6885 0.7967 0.2155
+   FET only 0.3766 0.4584 0.5111 0.0715 0.1894 0.3709 0.5891 0.7203 0.1624
+Table 12: Performance on the Legal Charge Retrieval (LCR) task with baselines. Highest results are in bold. “FET only” indicates using the four-element descriptions without case facts.
+To support reproducibility, we provide the full prompt templates used in our pipeline.
+Table 13 shows the prompt for charge prediction, and Table 14 presents the prompt used for generating four-element annotations in both BGE+FET(LLM) and BGE+FET(Expert, LLM).
+Prompt 1: Charge Prediction
+You are a legal expert specializing in criminal law.
+Based on the provided list of charges, determine which charges are applicable to the given case facts.
+Please note that you should only output the charge names, without any additional information.
+The charges must be selected from the provided list and should be separated by commas.
+[Crime List]
+[Case Facts]
+Table 13: Prompt used for charge prediction.
+Prompt 2: BGE+FET(LLM) and BGE+FET(Expert, LLM).
+You are a legal expert specializing in criminal law.
+Based on Chinese criminal law knowledge, analyze the following case facts and provide the following information in sequence: 1.
+The four elements of the crime:
+- Criminal Object: The tangible or intangible interests being infringed upon (e.g., personal rights such as life, or property rights such as money, vehicles).
+- Objective Aspect: The objective facts of the criminal activity, including key actions (e.g., theft, robbery) and consequences (e.g., injury, death, loss).
+- Criminal Subject: Typically general subjects; special subjects in certain crimes (e.g., government officials).
+- Subjective Aspect: Whether the act was intentional or negligent.
+2. Charge: Only output the specific crime name(s).
+3. Relevant Legal Articles: Only output the article number(s) of the relevant laws.
+Output format: JSON. For each crime involved in the case, provide a separate dictionary entry. [Output Sample]
+"Crime 1": {
+"Four elements": {
+"Criminal Object": "Personal rights: the victim Wang’s right to life; Property rights: vehicle.",
+"Objective Aspect": "The defendant Wu drove under the influence and collided with the victim Wang, causing Wang’s immediate death and vehicle damage.",
+"Criminal Subject": "Defendant Wu, the driver.",
+"Subjective Aspect": "Negligence"
+"Charge": "Traffic Accident Crime",
+"Relevant Legal Article": "Article 133"
+"Crime 2": {
+"Four elements": {
+"Criminal Object": "Social management order: infringement on the state’s document management system; Property rights: forged documents and related items.",
+"Objective Aspect": "Defendant 1 purchased equipment and materials to forge documents.
+Defendant 2 delivered the forged documents.
+Defendant 3 facilitated transactions via the internet, handling payments and document transfers.",
+"Criminal Subject": "Multiple defendants, all individuals with full criminal responsibility.",
+"Subjective Aspect": "Intentional"
+"Charge": "Forgery, Alteration, or Sale of Official Documents, Certificates, and Seals of State Organs",
+"Relevant Legal Article": "Article 280, Paragraph 1"
+Table 14: Prompt for generating four-element annotations used in FETLLM and FETExpert_Guided.
+Appendix H SCR results on the full LeCaRDv2 Dataset
 As presented in Table 15, we selected several representative methods based on sparse retrieval and dense retrieval for experiments on the full LeCaRDv2 dataset.
 All language models were not fine-tuned.
-The notes-guided FET method achieved the best performance among all language models, attaining top results in both R@500 and R@1000.
-The results indicate that the conclusions drawn from the full dataset are consistent with those from the subset, and the notes-guided method demonstrates strong performance.
+The expert-guided FET method achieved the best performance among all language models, attaining top results in both R@500 and R@1000.
+The results indicate that the conclusions drawn from the full dataset are consistent with those from the subset, and the expert-guided method demonstrates strong performance.
 Model R@100 R@200 R@500 R@1000
-BERT 0.1116 0.1493 0.2174 0.2819
+Legal-BERT 0.1116 0.1493 0.2174 0.2819
 Lawformer 0.2432 0.304 0.4054 0.4833
-ChatLaw 0.1045 0.1628 0.2791 0.3999
+ChatLaw-Text2Vec 0.1045 0.1628 0.2791 0.3999
 SAILER 0.2834 0.4033 0.6104 0.7568
 BGE 0.4085 0.5246 0.6855 0.7912
-FETLLM 0.4167 0.5388 0.7006 0.7925
-FETExpert_guided 0.4201 0.5396 0.7010 0.7927
+BGE+FET(GPT-4o) 0.4167 0.5388 0.7006 0.7925
+BGE+FET(Expert, GPT-4o) 0.4201 0.5396 0.7010 0.7927
 Table 15: SCR results on the full set of LeCaRDv2.
@@
 The expert-guided FET method achieved the best performance among all language models and attained the top results in both R@500 and R@1000.
-Appendix G A Case Study of LCR
+Appendix I A Case Study of LCR
 Table 16 presents a case study on the Crime of Embezzlement.
-By comparing the four elements annotated by experts for the crime in JUREX-4E, the case-specific four elements generated directly by the LLM, and those generated by the LLM with expert four elements of charge as guidance, we can observe that:
+By comparing the expert-FETs for the charges in JUREX-4E, the case-FETs generated directly by the LLM, and those generated by the LLM with expert-FETs of charge as guidance, we can observe that:
 1) Incorporating expert fine-grained annotations enables the model to better grasp the elements of a crime, thereby providing more precise element comparison.
@@
 Charge Crime of Embezzlement
-Experts Four-elements LLM-generated Experts-guided
-Four Elements
+Expert-FETs LLM-generated Expert-guided
+FETs
 Object: The ownership of property and integrity of duties in companies, enterprises, or other organizations.
@@
 Object: The ownership of property of companies, enterprises, or other organizations and the integrity of official duties.
-In this case, crime objetcs are the properties of the company, enterprise, or other organization where the perpetrator is employed.
-Objective Aspect: The act of using one’s position for personal gain by illegally appropriating the property of the organization, where the amount involved is significant.
+In this case, crime objects are the properties of the company, enterprise, or other organization where the perpetrator is employed.
+Objective Aspect: Using one’s position for personal gain by illegally appropriating the property of the organization, where the amount involved is significant.
 Specifically, B used his position to embezzle X yuan from the payments of six units for personal investment and consumption.
@@
 Subjective Aspect: Intentional, with the purpose of illegally possessing public or private property.
-Table 10: The results of the case four elements obtained through different methods in LCR.
-Notes refer to expert notes related to the charge retrieved during the search.
-LLM-generated and Experts-guided indicate whether using JUREX-4E’s four-elements of the crime to guide LLM in generating the four elements.
-Red parts mean the knowledge from JUREX-4E, while blue parts show the LLM’s internal knowledge.
-By incorporating JUREX-4E, the model better emphasizes conviction and sentencing related information and provides more detailed descriptions of critical case facts.
+Table 16: Results of the case-FETs using different methods for LCR.
+Expert-FETs refer to expert annotations of the predicted charges.
+LLM-generated and Expert-guided indicate whether the Expert-FETs of predicted charges are used to guide the LLM in generating the case-FETs.
+Red parts indicate knowledge from experts, while blue parts represent the LLM’s internal knowledge.
+By incorporating JUREX-4E, the model better emphasizes conviction- and sentencing-related information and provides more detailed descriptions of key case facts.